### v1

# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities in natural language processing tasks. However, their application in the biomedical domain presents unique challenges, particularly regarding factual accuracy and up-to-date knowledge integration. Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these challenges by combining the generative capabilities of LLMs with external knowledge retrieval. This comprehensive survey examines the application of RAG in the biomedical domain, focusing on its technological components, available datasets, and clinical applications. We present a systematic analysis of retrieval methods, ranking strategies, and generation models, while also exploring the challenges and future directions in this rapidly evolving field. Our work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems and identifies key areas for future research and development.
Keywords
Biomedical RAG, Large Language Model, Retrieval-Augmented Generation, Information Retrieval, Natural Language Processing, Biomedical NLP
Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 53, 81. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 219.
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 128. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 227. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 58. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 53, 81, 227.
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 230. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 131, 184.
Concurrent to our survey, several related surveys focused on biomedical-related LLMs. For example, Liu et al. 135 review medical LLMs and He et al. 81 discuss LLMs for healthcare. Similarly, Xiao 227 conducted a comprehensive overview of LLMs in medicine. Other discussed RAG systems in general domains, such as Fan et al. 53 and Gao et al. 65 focused on the architectures, training strategies, and application of RAG in general domains.
Our work differs from existing surveys in concentrating on the application of RAG in the biomedical domain, systematically reviewing its components, datasets, and challenges. Overall, this survey provides the following key contributions:
  1. 1.
    A systematic analysis of RAG technologies specifically adapted for biomedical applications, including specialized retrieval methods, ranking algorithms, and generation models.
  2. 2.
    A detailed examination of medical datasets, considering both knowledge source datasets and medical task datasets.
  3. 3.
    An exploration of current clinical applications and future directions, with particular attention to practical implementation challenges and emerging opportunities.
  4. 4.
    A structured framework for understanding the interplay between retrieval mechanisms, medical knowledge sources, and language models in biomedical RAG systems.
The remainder of this survey is organized as follows: Section 2 presents the background and the basic framework of biomedical RAG systems. Section 3 details the technical components, including retrieval methods, ranking algorithms, and generation models. Section 4 examines available datasets. Section 5 explores current clinical applications, and Section 6 discusses challenges and future directions.
Background
LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 247. Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 197. RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 65. This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 227.
In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources. The re-ranking method then sorts these references based on their relevance. Finally, the top-k most relevant references, combined with the original query, are processed by the LLM to generate the response. As Fig. 1 shows, the basic framework of biomedical RAG consists of four key components 53: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
Figure 1: Basic biomedical RAG framework.
RAG has been applied to a variety of biomedical applications. In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 141, enhances specialized decision-making through expert-guided LLMs 119, and facilitates rare disease diagnosis 240. This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 64 and advances genomic medicine through improved variant annotation and interpretation 142. In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 177. Across these applications, RAG leverages diverse data sources—including health records 105, 83, scientific literature 37, clinical guidelines 21, and genomic databases 142—to generate evidence-based responses that significantly improve both clinical care and research outcomes. The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
Technology
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 99, 243 and dense retrieval methods 168, 76, LLM-based retrieval 128, 163, knowledge graph enhancement techniques 108, 237, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 199, 177, 215. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 97, 94, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 165, 18, 73 and open-source models 7, 204, 156, 44 alongside domain adaptation techniques 58, 19 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
Retriever
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 65. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 218, 203. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains190. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this section, we describe various retrieval methodologies. We examine both standard retrieval approaches and how specific studies have modified these approaches for biomedical use, analyzing their respective strengths in handling domain-specific information structures and addressing unique retrieval challenges in the biomedical field.
Sparse Retrieval
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) 182 and TF-IDF (Term Frequency-Inverse Document Frequency) 106, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 1 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 42, 71 implementations to hybrid approaches 120, 13, 3, 115 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 109. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 234, 88, 92. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch 51, Whoosh 206, and Entrez API 56 also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 231, 120, 13, 3, 115, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
Table 1: Sparse retrieval methods in biomedical RAG systems
Method Name Retrieval Method Task Year
Clinfo.ai 141 Entrez API Medical QA 2023
MEDRAG toolkit 231 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS 88 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System 120 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA 13 BM25 and MedCPT Medical Question Answering 2024
CliniqIR 3 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery 115 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT 142 Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT 71 BM25 algorithm Medical QA 2025
Two-Layer RAG 42 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt 169 Entrez API for PubMed retrieval Scientific data analysis 2025
Dense Retrieval
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 246. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 190. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical information retrieval systems.
Type 1: PLMs PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 183, and Sentence Transformer 207 have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding 228, Dense Passage Retriever 111 and GIST Large Embedding 193 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText 26, RoBERTa 137, and SimCSE 63, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
Table 2: Classification of dense retriever types in biomedical RAG systems
Method Retriever Task Year
Type 1: PLMs
AskFDALabel 19 Sentence Transformer 207 FDA drug labeling extraction 2024
CaLM 168 BGE-embedding 228 with Chroma DB Supporting caregivers FM 2024
RAG-HPO 67 FastEmbed 205 Rare genetic disorders automated deep phenotyping 2024
RALL 76 Dense Passage Retriever 111 Lay language generation 2024
RT 129 BERT embeddings 46 Few-shot medical NER 2024
RAG-GPT 233 BGE-embedding 228 Breast cancer nursing care QA 2024
GNQA 41 HNSW graphs 147 Medical QA 2024
Klang et al. 116 GIST Large Embedding 193 with FAISS ICD-10-CM coding 2024
RAGPR 248 SBERT embeddings 183 Personalized physician recommendations 2025
JGCLLM 61 GLuCoSE-base-ja vectors 60 Genetic counseling support 2025
Type 2: Commercial APIs
DRAGON-AI 209 OpenAI Text-Embedding 164 with Chroma DB Ontology generation 2024
RISE 217 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG 121 UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT 130 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat 59 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT 139 Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT 198 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER 214 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP 20 Amazon Titan Text Embeddings v2 14 Nutrition-related question answering 2025
InfectA-Chat 186 Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG 107 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
Type 3: Domain adaptation PLMs
PM-Search 99 BioBERT 123 Clinical literature retrieval 2022
LADER 98 PubMedBERT 75 Biomedical literature retrieval 2023
CLEAR 140 BioBERT Clinical information extraction 2025
KG-RAG 194 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule 6 MedCPT, BioBERT, BioGPT 145 with Pinecone DB Biomedical QA 2024
Type 4: LLM-based embedding
BiomedRAG 128 MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM 163 GPT4All 16 with Chroma DB Surgical decision support 2024
Myers et al. 160 LLM2Vec (for comparison) 22 Clinical Information retrieval 2024
Type 5: Domain adaptation retriever
Self-BioRAG 93 MedCPT 97 Medical QA 2024
WeiseEule 6 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR 3 MedCPT and BM25 Diagnostic decision support 2024
MEDRAG 6 BM25, SPECTER 37, Contriever 92and MedCPT Biomedical RAG Tool 2024
RAMIE 242 MedCPT, Contriever and BMRetriever 234 Biomedical IR about dietary supplements 2025
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings dominate, alternatives are emerging. The RAP system 20 utilizes Amazon’s Titan Text Embeddings v2 14 for nutrition-related question answering. Other commercial APIs such as Anthropic’s Voyage-2 17 and Google’s Vertex AI embeddings 74, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 123, 75. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 146. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 146. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 75, BioBERT 123 and BioGPT 145, which have been widely adopted in biomedical RAG systems. For example, PM-Search 99 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER 98 for biomedical literature retrieval and in KG-RAG 194 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule 6 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 97 to achieve enhanced performance in biomedical QA.
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 247. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 128, which utilizes MedLLaMA 13b 224, SurgeryLLM 163, which employs GPT4All 16, and the comparative study by Myers et al. 160, which evaluates LLM2Vec 22 for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 138, Landmark Embedding 143, which provide promising dense retrieval alternatives in biomedical RAG.
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 92, MedCPT 97, SPECTER 37, BMRetriever 234, and the MEDRAG toolkit 128. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 93, 3, 242. For example, Jeong et al. 93 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework 6 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 242. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems 80. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, significantly enhancing the performance and response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we have compiled the most commonly used vector databases in Table 3. In Table LABEL:tab:Classification, it can be observed that the most popular vector stores are FAISS and Chroma, which were utilized 4 times and 3 times respectively.
Table 3: Common vector stores in biomedical RAG
Store Description
Chroma 36 AI-native open-source vector database with built-in functionality
Faiss 52 Library for efficient similarity search and clustering of dense vectors
Pinecone 172 Production-ready vector database for similarity search at scale
Weaviate 220 Graph-based vector search engine with semantic capabilities
HNSW 147 Hierarchical navigable small world graphs for approximate nearest neighbor search
Knowledge Graph Enhanced Retrieval
Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability171. Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source. As we can notice, the application areas vary significantly, including medical question answering 194, knowledge extraction 47, and diagnosis prediction 64.
Table 4: Recent biomedical RAG studies based on knowledge graphs
Method Task Knowledge Graph Used Year
KG-RAG 194 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE 108 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN 152 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle 237 Various medical text generation tasks UMLS 2024
DALK 125 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. 69 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 47 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO 229 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS 64 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT 151 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
As shown in Fig 2, the integration of knowledge graphs with LLMs in biomedical RAG applications can be analyzed through four key components:
  • Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. NetMe 2.0 47 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 229 mines PubMed literature for synthetic biology knowledge graphs. Differently, KG-RAG 194 utilizes the pre-existing SPOKE biomedical knowledge graph.
  • Query to knowledge graph transformation encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. ESCARGOT 151 employs Cypher queries for precise retrieval, KRAGEN 152 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 64 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
  • Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. DR.KNOWS 64 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 152 applies graph-of-thoughts for multi-hop relationships, and HEALIE 108 extracts paths incorporating both medical knowledge and socio-economic factors.
  • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency. KG-RAG 194 implements token-optimized retrieval to reduce consumption, ESCARGOT 151 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 47 transforms graph data into natural language prompts via its Graph-RAG module.
Figure 2: The Framework of Knowledge Graphs Based Biomedical RAG
Multi-modal Retrieval
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 4. As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 215, prescription label interpretation 208, lung cancer staging 211, digital pathology analysis 162, and automated radiology report generation 178, 199, 225, 226, 238.
Figure 3: Architectural workflow of multimodal biomedical RAG system
Table 5: Multimodal Biomedical RAG Systems
Model Task Modalities Year
Ranjit et al. 178 Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG 199 Radiology report generation X-Ray and text 2024
MMed-RAG 225 Medical VQA & Report generation Medical images and text 2024
MEAG 215 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al.  211 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG 162 Digital pathology QA Pathology images and text 2024
Raminedi et al. 177 Radiology report generation X-ray images and text 2024
RULE 226 Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. 208 Prescription label identification Image (labels) and text 2025
Hu et al. 87 Pathology report generation Whole slide images and text 2025
STREAM 238 Chest X-ray report generation Chest X-ray and text 2025
Figure 3 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 4. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 32 for images, specialized encoders for time-series data 32); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.178, FactMM-RAG 199 and STREAM 238 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG 225 and RULE 226 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG 215 system uses specialized encoders for eye movement data alongside text, Raminedi et al. 162 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. 208 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
Reranker
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 190. Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 3 and MEDRAG 6 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 209 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 6 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 94 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 5 utilizes a specialized cross-encoder model known as BGE M3-Embedding 31 to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 97 trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 41 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 39 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai 141 combines LLM classification with BM25 to re-rank retrieved articles.
Generation
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 122. This section examines various generation models employed in recent biomedical RAG studies.
Current generation models employed in biomedical RAG systems can be categorized into three distinct classes based on their development approach and specialization: general-domain open-source models, commercial proprietary models, and biomedical-specialized models. These categories present different trade-offs in terms of accessibility, customization potential, and domain-specific performance. Table LABEL:tab:llm_taxonomy showcases representative models from each category and highlights recent studies that utilize these models. It is important to note that our analysis of biomedical-specialized LLMs focuses specifically on models trained directly on biomedical or healthcare corpora, deliberately excluding applications developed through prompt engineering or Chain-of-Thought (CoT) methodologies.
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 247. Among open-source LLMs, LLaMa 2 210 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 7, Mistral 7B 155, and T5 176 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 156 and Gemma 3 72. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 217, 84, 119, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies 166, 242, 11 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
Table 6: Taxonomic Classification of LLMs for Embedding Applications
Model Parameters Architecture Studies
Open-Source LLMs
T5 176 0.06-11B Encoder-Decoder VAIV 115, CLEAR 140, DR.KNOWS 64
ChatGLM3 70 6B Decoder-only COPD 244, ChatZOC 144
LLaMA 2 210 7-70B Decoder-only Guo et al. 76, KREIMEYER et al. 117, CaLM 168, MMRAG 241, Yu et al. 239, RAMIE 242, EyeGPT 33, KG-RAG 194, JMLR 219, BiomedRAG 128, Self-BioRAG 93, CaLM 168, Du et al. 49, SurgeryLLM 163, Kreimeyer et al. 117
LLaMA 3 7 8-70B Decoder-only MMRAG 241, Fatharanihttps et al. 54, RAMIE 242, Woo et al. 222, RAG-HPO 67, i-MedRAG 232, Hewitt et al. 84, PodGPT 94, SurgeryLLM 163
Phi-3 Mini 154 3.8B Decoder-only Fatharanihttps et al. 54
Mistral 7B 155 7B Decoder-only Boulos et al. 110, RAMIE 242, Woo et al. 222, RAGPR 248, LITURAt 169, Kreimeyer et al. 117
QWen 175 32B Decoder-only NEKO 229, Thetbanthad 208
Baichuan 216 7-13B Decoder-only ChatZOC 144
Falcon 12 7-180B Decoder-only AskFDALabel 19, CaLM 168
Zephyr 213 7B Decoder-only Moser et al. 159
AceGPT 89 7-13B Decoder-only InfectA-Chat 186
Gemma 62 1-27B Decoder-only PodGPT 94
Deepseek R1 44 671B Decoder-only Feng et al. 55
Commercial LLMs
ChatGPT-3.5/4/4o 165 Proprietary Decoder-only Woo et al. 222, RAGPR 248, ChatZOC 144, KG-RAG 194, Clinfo.ai 141, Yu et al. 239, ChatENT 139, CaLM 168, Rau et al. 180, Endo-chat 59, BiomedRAG 128, Lu 142, Tan 15, GNQA 40, Puts 173, DRAGON-AI 209, Choi 35, RISE 217, Du et al. 49,i-MedRAG 232 , Hewitt et al. 84, RECTIFIER 214, Kainer et al. 107, Gong et al. 71, VAIA 115, RareDxGPT 240, Selcuk et al. 186, FAVOR-GPT 130, Markey et al. 148,
Claude-3.5 18 Proprietary Decoder-only Woo et al. 222, RISE 217, Hewitt et al. 84
Gemini 73 Proprietary Multimodal Upadhyaya et al. 215, Tozuka et al. 211, MEREDITH 119
Biomedical Specialized LLMs
MedAlpaca 79 7-13B LLaMA-based RAMIE 242
PMC-LLaMA 223 7B LLaMA-based RAMIE 242, BiomedRAG 128, Ozaki et al. 166
BioMistral 118 7B Mistral-based RAMIE 242, pRAGe 28
MEDITRON 34 7-70B LLaMA-based Ozaki et al. 166, Alkaeed et al. 11
Meerkat 113 7B Mistral-based Sohn et al. 192
Domain Adaptation
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 189. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 92 and MedCPT 97) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 234) and domain-specific architectural modifications (exemplified by LADER 98 with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 97) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 128). Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, significantly enhance the system’s capacity to prioritize documents based on domain-relevant criteria. Furthermore, they facilitate the identification of subtle domain-specific signals that general models typically fail to detect, thereby improving retrieval precision for specialized medical queries.
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 58, SeRTS 88), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 19), and continued pre-training (RALL 76, ChatENT 139) enables more accurate synthesis and explication of biomedical information. JLMR 219 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns essential for healthcare and life sciences applications. Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, ultimately advancing the state-of-the-art in biomedical information retrieval and synthesis.
Datasets
The evaluation and development of biomedical RAG systems rely heavily on appropriate datasets. In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task datasets. This categorization reflects the dual nature of RAG systems, which require both comprehensive knowledge bases and task-specific evaluation benchmarks.
Knowledge Source Datasets
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 66, 53. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 90, 139, 93.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries 6, 234.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS 25, PMC, and Medical textbooks 95 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies 35, 49, 248 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 120, 150, 233, 20, 89, 187 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
Table 7: Comprehensive Medical Knowledge Sources
Source Type Description Studies
MIMIC-IV 103 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 159
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 141, MEDRAG 231, Self-BioRAG 93, JMLR 219, Ascle 237, RISE 217, MEREDITH 119, LITURAt 169, DR.KNOWS 64, Aftab 6, CliniqIR 3, PM-Search 99, Kreimeyer 117, Kim 115
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 93, Soong et al. 196, PodGPT 94
TripClick 181 Literature Biomedical literature search and retrieval dataset LADER 98
GNQA 40 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 40
UMLS 25 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 128, CLEAR 140, Guo 76, CliniqIR 3
ICD-10 30 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 173
HPO 102 Knowledge Base Human phenotype ontology Albayrak et al. 9, RAG-HPO 67, Kainer 107
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 234
AlzKB 151 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 151, KRAGEN 152
DrugBank 221 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 234, Kim 115
SPOKE 158 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 194
COD 144 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 144
HEALIE KG108 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 108
ADRD Caregiving 168 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 168
StatPearls 21 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 231, i-MedRAG 232
Medical Textbooks 95 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 231, Self-BioRAG 93, JMLR 219, i-MedRAG 232
Ophthalmology textbooks 33 Educational 14 specialized ophthalmology textbooks EyeGPT 33
CheXpert 91 Image-Report 224,316 chest radiographs with associated reports CLEAR 140
Medical Task Datasets
Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 81, 135. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups 61, 198, 217 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
Table 8: Comprehensive Biomedical Tasks Datasets
Dataset Year Task Studies
ChemProt 202 2010 Information Extraction BiomedRAG 128, ChemProt
DDI 185 2013 Information Extraction BiomedRAG 128, ChemProt
NCBI 48 2014 Entity Recognition GeneGPT 100, NetMe 47, RT 129
BioASQ 212 2015 Information Retrieval BMRETRIEVER 234, SeRTS 88
MIMIC-III 105 2016 Text Summarization DR.KNOWS 64, CliniqIR 3
BC5CDR 126 2016 Entity Recognition RT 129
DC3 50 2019 Diagnostic classification CliniqIR 3
MIMIC-CXR 104 2019 Text Summarization Ascle 237
MedQuAD 23 2019 QA Ascle 237
PubMedQA 96 2019 Multiple-choice BMRETRIEVER 234, Liévin et al. 133, PodGPT 94
MMLU 82 2021 Multiple-choice Self-BioRAG 93, JMLR 219, i-MedRAG 232, InfectA-Chat 186, PodGPT 94
MedQA 95 2021 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, i-MedRAG 232, PodGPT 94
HealthSearchQA 188 2022 QA
MedMCQA 167 2022 Multiple-choice Self-BioRAG 93, JMLR 219, Ascle 237, Liévin et al. 133, EyeGPT 33, PodGPT 94
ChatDoctor 132 2023 Dialogue EyeGPT 33
MedAlpaca 79 2023 Dialogue EyeGPT 33
GeneTuring 86 2023 Genomics QA GeneGPT 100
GIT 127 2023 Information Extraction BiomedRAG 128
MIRAGE 231 2024 Multiple-choice & QA MEDRAG 231
MedExpQA 13 2024 Multilingual Medical QA MedExpQA 13, PodGPT 94
PubMedRS-200 141 2024 Medical QA Clinifo.ai 141
MedicineQA 90 2024 Multi-round dialogue RagPULSE 90
CELLS 76 2024 Lay language generation Guo 76
Applications
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 136. In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
Clinical Decision Support Systems
Medical Question Answering
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 194. This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai 141 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL 76 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information. Self-BioRAG 93 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 6 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
Diagnostic and Treatment Decision Support
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 201. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs 119 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS 64 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR 3 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 163 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER 214 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
Rare Disease Identification and Management
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 149. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. 240 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO 67 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. 236 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. Some authors 114 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
Clinical Report Generation
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 191. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 78. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 10.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 230. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 179, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors 8 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines 24 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG 195 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors 200 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
Precision Medicine Applications
Genomic Medicine
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 29.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT 130 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. 142 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO 67 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 61 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 142 and phenotypic and genotypic analysis 67, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
Personalized Health Management
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 174. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE 108 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE 217 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System 120 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP 20 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
Healthcare Education
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 124. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT 139 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT 33 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. 187 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval. In nursing education, Zhao et al 245 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.
Clinical Research
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. 68.
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT 94 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt 169 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
Challenges and Future Direction
Trustworthiness
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 53, 45. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 66. Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 122. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora 81, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
Multilinguality
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 161, 101. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 157, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
Multimodality
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray 104, 170, 134), time series (e.g. ECG 239), sound recordings (e.g. lung sounds recorded 57), and video 77. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 249. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the Biomedical domain and improving its accuracy and trustworthiness in the future.
Restricted Computing Resources
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 53. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 43.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 27, 85. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts 235. Models such as QWQ 175 and Mistral Small 3.1 156 demonstrate competitive performance despite having significantly reduced parameter counts. Applying Portable LLMs 112 as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
Privacy
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 153. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 38 and the Health Insurance Portability and Accountability Act 2. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.
Conclusion
In conclusion, this comprehensive survey has provided a detailed overview of the application of RAG in the biomedical domain. It has examined the basic framework of biomedical RAG and discussed its retriever, reranker, and generation modules, as well as explored KG-enhanced biomedical RAG and multi-modal biomedical RAG. The survey has also explored clinical applications where biomedical RAG can be applied and reviewed recent studies in this field. Additionally, the paper has identified key challenges, including trustworthiness, multilingual capabilities, multimodal integration, limited computational resources, and privacy concerns. Overall, this work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems, ultimately contributing to the development of more accurate, accessible, and impactful healthcare AI systems that can meaningfully enhance patient care through evidence-based information access and support.