### v2

# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities in natural language processing tasks. However, their application in the biomedical domain presents unique challenges, particularly regarding factual accuracy and up-to-date knowledge integration. Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these challenges by combining the generative capabilities of LLMs with external knowledge retrieval. This comprehensive survey examines the application of RAG in the biomedical domain, focusing on its technological components, available datasets, and clinical applications. We present a systematic analysis of retrieval methods, ranking strategies, and generation models, while also exploring the challenges and future directions in this rapidly evolving field. Our work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems and identifies key areas for future research and development.
Keywords
Biomedical RAG, Large Language Model, Retrieval-Augmented Generation, Information Retrieval, Natural Language Processing, Biomedical NLP
Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks 1, 2. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedical domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes 3.
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination 4. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge 5. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges 6. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) enhance the accuracy and reliability of medical information processing; and (4) reduce hallucination by grounding responses in the verified medical literature 1, 2, 5.
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability 7. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 8, 9.
Concurrent to our survey, several related surveys focused on biomedical-related LLMs. For example, Liu et al. 10 review medical LLMs and He et al. 2 discuss LLMs for healthcare. Similarly, Xiao 5 conducted a comprehensive overview of LLMs in medicine. Other discussed RAG systems in general domains, such as Fan et al. 1 and Gao et al. 11 focused on the architectures, training strategies, and application of RAG in general domains.
Our work differs from existing surveys in concentrating on the application of RAG in the biomedical domain, systematically reviewing its components, datasets, and challenges. Overall, this survey provides the following key contributions:
  1. 1.
    A systematic analysis of RAG technologies specifically adapted for biomedical applications, including specialized retrieval methods, ranking algorithms, and generation models.
  2. 2.
    A detailed examination of medical datasets, considering both knowledge source datasets and medical task datasets.
  3. 3.
    An exploration of current clinical applications and future directions, with particular attention to practical implementation challenges and emerging opportunities.
  4. 4.
    A structured framework for understanding the interplay between retrieval mechanisms, medical knowledge sources, and language models in biomedical RAG systems.
The remainder of this survey is organized as follows: Section 2 presents the background and the basic framework of biomedical RAG systems. Section 3 details the technical components, including retrieval methods, ranking algorithms, and generation models. Section 4 examines available datasets. Section 5 explores current clinical applications, and Section 6 discusses challenges and future directions.
Background
LLMs have disrupted NLP tasks with their ability to generate human-like text across diverse tasks 12. Despite their impressive capabilities, LLMs often struggle with knowledge limitations, hallucinations, and outdated information, particularly in specialized domains requiring precise information 13. RAG addresses these limitations by dynamically retrieving relevant external knowledge to supplement the LLM’s parametric knowledge before generating responses 11. This approach has shown particular promise in domain-specific applications, such as biomedicine, where accuracy, recency, and specialized knowledge are critical for clinical decision support, research analysis, and patient care 5.
In a typical workflow, a query first passes through the retriever, which searches relevant documents from knowledge sources. The re-ranking method then sorts these references based on their relevance. Finally, the top-k most relevant references, combined with the original query, are processed by the LLM to generate the response. As Fig. 1 shows, the basic framework of biomedical RAG consists of four key components 1: (1) Retriever: Responsible for indexing and retrieving relevant documents from various knowledge sources; Given a query $q$ and a collection of knowledge sources $K$, a retriever function $R(q,K)$ outputs a set of relevant documents $D=\{d_{1},d_{2},\ldots,d_{n}\}$, where each $d_{i}\in K$; (2) Knowledge Source: $K$ represents a heterogeneous collection of structured and unstructured data, including but not limited to medical literature, clinical guidelines, and databases; (3) Re-ranking Method:Re-sorting and prioritizing retrieved references, A re-ranking function $S(q,D)$ takes the query $q$ and the retrieved document set $D$ as input and outputs a re-ranked list of documents $D^{\prime}=[d_{1}^{\prime},d_{2}^{\prime},\ldots,d_{m}^{\prime}]$, where $D^{\prime}\subseteq D$ and the order reflects relevance to $q$; and (4) LLM: A LLM $G(q,D^{\prime})$ takes the original query $q$ and the re-ranked retrieved context $D^{\prime}$ as input to generate a final response $y$.
Figure 1: Basic biomedical RAG framework.
RAG has been applied to a variety of biomedical applications. In clinical decision support systems, RAG powers evidence-based question-answering platforms like Clinfo.ai 14, enhances specialized decision-making through expert-guided LLMs 15, and facilitates rare disease diagnosis 16. This technology also strengthens diagnostic prediction by connecting electronic health records (EHRs) with medical knowledge repositories 17 and advances genomic medicine through improved variant annotation and interpretation 18. In the realm of medical imaging, RAG enhances report generation by seamlessly incorporating relevant prior studies and clinical guidelines 19. Across these applications, RAG leverages diverse data sources—including health records 20, 21, scientific literature 22, clinical guidelines 23, and genomic databases 18—to generate evidence-based responses that significantly improve both clinical care and research outcomes. The growing body of evidence consistently demonstrates that RAG holds remarkable promise for advancing information management and decision support in the biomedical and healthcare domains.
Technology
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse 24, 25 and dense retrieval methods 26, 27, LLM-based retrieval 4, 28, knowledge graph enhancement techniques 29, 30, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts 31, 19, 32. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents 33, 34, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source 35, 36, 37 and open-source models 38, 39, 40, 41 alongside domain adaptation techniques 6, 42 that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
Retriever
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries 11. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 43, 44. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domains45. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this section, we describe various retrieval methodologies. We examine both standard retrieval approaches and how specific studies have modified these approaches for biomedical use, analyzing their respective strengths in handling domain-specific information structures and addressing unique retrieval challenges in the biomedical field.
Sparse Retrieval
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) 46 and TF-IDF (Term Frequency-Inverse Document Frequency) 47, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 1 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 48, 49 implementations to hybrid approaches 50, 51, 52, 53 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization 54. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements 55, 56, 57. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch 58, Whoosh 59, and Entrez API 60 also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques 61, 50, 51, 52, 53, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
Table 1: Sparse retrieval methods in biomedical RAG systems
Method Name Retrieval Method Task Year
Clinfo.ai 14 Entrez API Medical QA 2023
MEDRAG toolkit 61 Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS 56 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System 50 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA 51 BM25 and MedCPT Medical Question Answering 2024
CliniqIR 52 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery 53 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT 18 Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT 49 BM25 algorithm Medical QA 2025
Two-Layer RAG 48 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt 62 Entrez API for PubMed retrieval Scientific data analysis 2025
Dense Retrieval
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 63. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching 45. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical information retrieval systems.
Type 1: PLMs PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT 64, and Sentence Transformer 65 have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding 66, Dense Passage Retriever 67 and GIST Large Embedding 68 have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText 69, RoBERTa 70, and SimCSE 71, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
Table 2: Classification of dense retriever types in biomedical RAG systems
Method Retriever Task Year
Type 1: PLMs
AskFDALabel 42 Sentence Transformer 65 FDA drug labeling extraction 2024
CaLM 26 BGE-embedding 66 with Chroma DB Supporting caregivers FM 2024
RAG-HPO 72 FastEmbed 73 Rare genetic disorders automated deep phenotyping 2024
RALL 27 Dense Passage Retriever 67 Lay language generation 2024
RT 74 BERT embeddings 75 Few-shot medical NER 2024
RAG-GPT 76 BGE-embedding 66 Breast cancer nursing care QA 2024
GNQA 77 HNSW graphs 78 Medical QA 2024
Klang et al. 79 GIST Large Embedding 68 with FAISS ICD-10-CM coding 2024
RAGPR 80 SBERT embeddings 64 Personalized physician recommendations 2025
JGCLLM 81 GLuCoSE-base-ja vectors 82 Genetic counseling support 2025
Type 2: Commercial APIs
DRAGON-AI 83 OpenAI Text-Embedding 84 with Chroma DB Ontology generation 2024
RISE 85 OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG 86 UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT 87 Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat 88 Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT 89 Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT 90 Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER 91 Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP 92 Amazon Titan Text Embeddings v2 93 Nutrition-related question answering 2025
InfectA-Chat 94 Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG 95 OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
Type 3: Domain adaptation PLMs
PM-Search 24 BioBERT 96 Clinical literature retrieval 2022
LADER 97 PubMedBERT 98 Biomedical literature retrieval 2023
CLEAR 99 BioBERT Clinical information extraction 2025
KG-RAG 100 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule 101 MedCPT, BioBERT, BioGPT 102 with Pinecone DB Biomedical QA 2024
Type 4: LLM-based embedding
BiomedRAG 4 MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM 28 GPT4All 103 with Chroma DB Surgical decision support 2024
Myers et al. 104 LLM2Vec (for comparison) 105 Clinical Information retrieval 2024
Type 5: Domain adaptation retriever
Self-BioRAG 106 MedCPT 33 Medical QA 2024
WeiseEule 101 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR 52 MedCPT and BM25 Diagnostic decision support 2024
MEDRAG 101 BM25, SPECTER 22, Contriever 57and MedCPT Biomedical RAG Tool 2024
RAMIE 107 MedCPT, Contriever and BMRetriever 55 Biomedical IR about dietary supplements 2025
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings dominate, alternatives are emerging. The RAP system 92 utilizes Amazon’s Titan Text Embeddings v2 93 for nutrition-related question answering. Other commercial APIs such as Anthropic’s Voyage-2 108 and Google’s Vertex AI embeddings 109, though not specifically represented in Table LABEL:tab:Classification, are alternatives for biomedical RAG applications.
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 96, 98. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature 110. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpus, these models develop a more nuanced understanding of biomedical language, resulting in significantly improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts 110. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT 98, BioBERT 96 and BioGPT 102, which have been widely adopted in biomedical RAG systems. For example, PM-Search 24 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER 97 for biomedical literature retrieval and in KG-RAG 100 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule 101 combine multiple biomedical models including BioBERT, BioGPT, and MedCPT 33 to achieve enhanced performance in biomedical QA.
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text 12. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings that outperform traditional embedding approaches. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG 4, which utilizes MedLLaMA 13b 111, SurgeryLLM 28, which employs GPT4All 103, and the comparative study by Myers et al. 104, which evaluates LLM2Vec 105 for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec 112, Landmark Embedding 113, which provide promising dense retrieval alternatives in biomedical RAG.
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever 57, MedCPT 33, SPECTER 22, BMRetriever 55, and the MEDRAG toolkit 4. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies 106, 52, 107. For example, Jeong et al. 106 proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework 101 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIE utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments 107. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems 114. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, significantly enhancing the performance and response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we have compiled the most commonly used vector databases in Table 3. In Table LABEL:tab:Classification, it can be observed that the most popular vector stores are FAISS and Chroma, which were utilized 4 times and 3 times respectively.
Table 3: Common vector stores in biomedical RAG
Store Description
Chroma 115 AI-native open-source vector database with built-in functionality
Faiss 116 Library for efficient similarity search and clustering of dense vectors
Pinecone 117 Production-ready vector database for similarity search at scale
Weaviate 118 Graph-based vector search engine with semantic capabilities
HNSW 78 Hierarchical navigable small world graphs for approximate nearest neighbor search
Knowledge Graph Enhanced Retrieval
Knowledge graph retrievers offer significant advantages in biomedical RAG by providing structured, interlinked domain knowledge that overcomes several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainability119. Table 4 presents knowledge graph-based biomedical RAG studies published between 2024 and 2025, outlining each system’s application domains, and knowledge graph source. As we can notice, the application areas vary significantly, including medical question answering 100, knowledge extraction 120, and diagnosis prediction 17.
Table 4: Recent biomedical RAG studies based on knowledge graphs
Method Task Knowledge Graph Used Year
KG-RAG 100 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE 29 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN 121 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle 30 Various medical text generation tasks UMLS 2024
DALK 122 Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. 123 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 120 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO 124 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS 17 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT 125 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
As shown in Fig 2, the integration of knowledge graphs with LLMs in biomedical RAG applications can be analyzed through four key components:
Figure 2: The Framework of Knowledge Graphs Based Biomedical RAG
  • Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. NetMe 2.0 120 constructs biomedical knowledge graphs on-the-fly with OntoTagMe for entity linking while NEKO 124 mines PubMed literature for synthetic biology knowledge graphs. Differently, KG-RAG 100 utilizes the pre-existing SPOKE biomedical knowledge graph.
  • Query structuration encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. ESCARGOT 125 employs Cypher queries for precise retrieval, KRAGEN 121 vectorizes knowledge graphs for semantic searching, and DR.KNOWS 17 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
  • Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. DR.KNOWS 17 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN 121 applies graph-of-thoughts for multi-hop relationships, and HEALIE 29 extracts paths incorporating both medical knowledge and socio-economic factors.
  • Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts for LLMs in ways that maximize factual accuracy while optimizing token usage and computational efficiency. KG-RAG 100 implements token-optimized retrieval to reduce consumption, ESCARGOT 125 selectively integrates knowledge through dynamic Graph of Thoughts, and NetMe 2.0 120 transforms graph data into natural language prompts via its Graph-RAG module.
Multi-modal Retrieval
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches 126. As shown in Table 5, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection 32, prescription label interpretation 127, lung cancer staging 128, digital pathology analysis 129, and automated radiology report generation 130, 31, 131, 132, 133.
Figure 3: Architectural workflow of multimodal biomedical RAG system
Table 5: Multimodal Biomedical RAG Systems
Model Task Modalities Year
Ranjit et al. 130 Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG 31 Radiology report generation X-Ray and text 2024
MMed-RAG 131 Medical VQA & Report generation Medical images and text 2024
MEAG 32 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al.  128 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG 129 Digital pathology QA Pathology images and text 2024
Raminedi et al. 19 Radiology report generation X-ray images and text 2024
RULE 132 Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. 127 Prescription label identification Image (labels) and text 2025
Hu et al. 134 Pathology report generation Whole slide images and text 2025
STREAM 133 Chest X-ray report generation Chest X-ray and text 2025
Figure 3 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs 126. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers 135 for images, specialized encoders for time-series data 135); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.130, FactMM-RAG 31 and STREAM 133 use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG 131 and RULE 132 build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG 32 system uses specialized encoders for eye movement data alongside text, Raminedi et al. 129 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. 127 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
Reranker
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models 45. Positioned between retrievers and generators, rerankers apply sophisticated algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR 52 and MEDRAG 101 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI 83 employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. Content-based rerankers, in contrast, focus on content similarity; notably, WeiseEule 101 introduces a keyword frequency-based ranking system specifically designed for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT 34 uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system 136 utilizes a specialized cross-encoder model known as BGE M3-Embedding 137 to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT 33 trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA 77 employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC 138 implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai 14 combines LLM classification with BM25 to re-rank retrieved articles.
Generation
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions 139. This section examines various generation models employed in recent biomedical RAG studies.
Current generation models employed in biomedical RAG systems can be categorized into three distinct classes based on their development approach and specialization: general-domain open-source models, commercial proprietary models, and biomedical-specialized models. These categories present different trade-offs in terms of accessibility, customization potential, and domain-specific performance. Table LABEL:tab:llm_taxonomy showcases representative models from each category and highlights recent studies that utilize these models. It is important to note that our analysis of biomedical-specialized LLMs focuses specifically on models trained directly on biomedical or healthcare corpora, deliberately excluding applications developed through prompt engineering or Chain-of-Thought (CoT) methodologies.
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency 12. Among open-source LLMs, LLaMa 2 140 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 38, Mistral 7B 141, and T5 142 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 40 and Gemma 3 143. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications 85, 144, 15, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies 145, 107, 146 have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
Table 6: Taxonomic Classification of LLMs for Embedding Applications
Model Parameters Architecture Studies
Open-Source LLMs
T5 142 0.06-11B Encoder-Decoder VAIV 53, CLEAR 99, DR.KNOWS 17
ChatGLM3 147 6B Decoder-only COPD 148, ChatZOC 149
LLaMA 2 140 7-70B Decoder-only Guo et al. 27, KREIMEYER et al. 150, CaLM 26, MMRAG 151, Yu et al. 152, RAMIE 107, EyeGPT 153, KG-RAG 100, JMLR 3, BiomedRAG 4, Self-BioRAG 106, CaLM 26, Du et al. 154, SurgeryLLM 28, Kreimeyer et al. 150
LLaMA 3 38 8-70B Decoder-only MMRAG 151, Fatharanihttps et al. 155, RAMIE 107, Woo et al. 156, RAG-HPO 72, i-MedRAG 157, Hewitt et al. 144, PodGPT 34, SurgeryLLM 28
Phi-3 Mini 158 3.8B Decoder-only Fatharanihttps et al. 155
Mistral 7B 141 7B Decoder-only Boulos et al. 159, RAMIE 107, Woo et al. 156, RAGPR 80, LITURAt 62, Kreimeyer et al. 150
QWen 160 32B Decoder-only NEKO 124, Thetbanthad 127
Baichuan 161 7-13B Decoder-only ChatZOC 149
Falcon 162 7-180B Decoder-only AskFDALabel 42, CaLM 26
Zephyr 163 7B Decoder-only Moser et al. 164
AceGPT 165 7-13B Decoder-only InfectA-Chat 94
Gemma 166 1-27B Decoder-only PodGPT 34
Deepseek R1 41 671B Decoder-only Feng et al. 167
Commercial LLMs
ChatGPT-3.5/4/4o 35 Proprietary Decoder-only Woo et al. 156, RAGPR 80, ChatZOC 149, KG-RAG 100, Clinfo.ai 14, Yu et al. 152, ChatENT 89, CaLM 26, Rau et al. 168, Endo-chat 88, BiomedRAG 4, Lu 18, Tan 169, GNQA 77, Puts 170, DRAGON-AI 83, Choi 171, RISE 85, Du et al. 154,i-MedRAG 157 , Hewitt et al. 144, RECTIFIER 91, Kainer et al. 95, Gong et al. 49, VAIA 53, RareDxGPT 16, Selcuk et al. 94, FAVOR-GPT 87, Markey et al. 172,
Claude-3.5 36 Proprietary Decoder-only Woo et al. 156, RISE 85, Hewitt et al. 144
Gemini 37 Proprietary Multimodal Upadhyaya et al. 32, Tozuka et al. 128, MEREDITH 15
Biomedical Specialized LLMs
MedAlpaca 173 7-13B LLaMA-based RAMIE 107
PMC-LLaMA 174 7B LLaMA-based RAMIE 107, BiomedRAG 4, Ozaki et al. 145
BioMistral 175 7B Mistral-based RAMIE 107, pRAGe 176
MEDITRON 177 7-70B LLaMA-based Ozaki et al. 145, Alkaeed et al. 146
Meerkat 178 7B Mistral-based Sohn et al. 179
Domain Adaptation
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications 180. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever 57 and MedCPT 33) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER 55) and domain-specific architectural modifications (exemplified by LADER 97 with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker 33) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer 4). Notably, the aforementioned Model-based rerankers and LLM-based rerankers (Sec Reranker) can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, significantly enhance the system’s capacity to prioritize documents based on domain-relevant criteria. Furthermore, they facilitate the identification of subtle domain-specific signals that general models typically fail to detect, thereby improving retrieval precision for specialized medical queries.
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE 6, SeRTS 56), parameter-efficient approaches like Low-Rank Adaptation (LoRA) (implemented in AskFDALabel 42), and continued pre-training (RALL 27, ChatENT 89) enables more accurate synthesis and explication of biomedical information. JLMR 3 joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns essential for healthcare and life sciences applications. Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, ultimately advancing the state-of-the-art in biomedical information retrieval and synthesis.
Datasets
The evaluation and development of biomedical RAG systems rely heavily on appropriate datasets. In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task datasets. This categorization reflects the dual nature of RAG systems, which require both comprehensive knowledge bases and task-specific evaluation benchmarks.
Knowledge Source Datasets
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources 181, 1. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices 182, 89, 106.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries 101, 55.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS 183, PMC, and Medical textbooks 184 demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies 171, 154, 80 have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies 50, 185, 76, 92, 165, 186 focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
Table 7: Comprehensive Medical Knowledge Sources
Source Type Description Studies
MIMIC-IV 187 EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. 164
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai 14, MEDRAG 61, Self-BioRAG 106, JMLR 3, Ascle 30, RISE 85, MEREDITH 15, LITURAt 62, DR.KNOWS 17, Aftab 101, CliniqIR 52, PM-Search 24, Kreimeyer 150, Kim 53
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG 106, Soong et al. 188, PodGPT 34
TripClick 189 Literature Biomedical literature search and retrieval dataset LADER 97
GNQA 77 Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA 77
UMLS 183 Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG 4, CLEAR 99, Guo 27, CliniqIR 52
ICD-10 190 Knowledge Base International statistical classification of diseases and related health problems Puts et al. 170
HPO 191 Knowledge Base Human phenotype ontology Albayrak et al. 192, RAG-HPO 72, Kainer 95
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER 55
AlzKB 125 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT 125, KRAGEN 121
DrugBank 193 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER 55, Kim 53
SPOKE 194 Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG 100
COD 149 Knowledge Base Comprehensive ophthalmic dataset ChatZOC 149
HEALIE KG29 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. 29
ADRD Caregiving 26 Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM 26
StatPearls 23 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG 61, i-MedRAG 157
Medical Textbooks 184 Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG 61, Self-BioRAG 106, JMLR 3, i-MedRAG 157
Ophthalmology textbooks 153 Educational 14 specialized ophthalmology textbooks EyeGPT 153
CheXpert 195 Image-Report 224,316 chest radiographs with associated reports CLEAR 99
Medical Task Datasets
Medical task datasets represent essential benchmarking instruments for evaluating RAG systems within healthcare applications 2, 10. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups 81, 90, 85 in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
Table 8: Comprehensive Biomedical Tasks Datasets
Dataset Year Task Studies
ChemProt 196 2010 Information Extraction BiomedRAG 4, ChemProt
DDI 197 2013 Information Extraction BiomedRAG 4, ChemProt
NCBI 198 2014 Entity Recognition GeneGPT 199, NetMe 120, RT 74
BioASQ 200 2015 Information Retrieval BMRETRIEVER 55, SeRTS 56
MIMIC-III 20 2016 Text Summarization DR.KNOWS 17, CliniqIR 52
BC5CDR 201 2016 Entity Recognition RT 74
DC3 202 2019 Diagnostic classification CliniqIR 52
MIMIC-CXR 203 2019 Text Summarization Ascle 30
MedQuAD 204 2019 QA Ascle 30
PubMedQA 205 2019 Multiple-choice BMRETRIEVER 55, Liévin et al. 206, PodGPT 34
MMLU 207 2021 Multiple-choice Self-BioRAG 106, JMLR 3, i-MedRAG 157, InfectA-Chat 94, PodGPT 34
MedQA 184 2021 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, i-MedRAG 157, PodGPT 34
HealthSearchQA 208 2022 QA
MedMCQA 209 2022 Multiple-choice Self-BioRAG 106, JMLR 3, Ascle 30, Liévin et al. 206, EyeGPT 153, PodGPT 34
ChatDoctor 210 2023 Dialogue EyeGPT 153
MedAlpaca 173 2023 Dialogue EyeGPT 153
GeneTuring 211 2023 Genomics QA GeneGPT 199
GIT 212 2023 Information Extraction BiomedRAG 4
MIRAGE 61 2024 Multiple-choice & QA MEDRAG 61
MedExpQA 51 2024 Multilingual Medical QA MedExpQA 51, PodGPT 34
PubMedRS-200 14 2024 Medical QA Clinifo.ai 14
MedicineQA 182 2024 Multi-round dialogue RagPULSE 182
CELLS 27 2024 Lay language generation Guo 27
Applications
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome 213. In this section, we analyze key healthcare domains where biomedical RAG systems have demonstrated substantial utility and are organized according to their functional contributions to clinical practice, precision medicine, and healthcare knowledge dissemination.
Clinical Decision Support Systems
Medical Question Answering
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need 100. This application is essential because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several sophisticated medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai 14 provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL 27 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the critical need for accessible medical information. Self-BioRAG 106 implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule 101 integrates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
Diagnostic and Treatment Decision Support
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis 214. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs 15 support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS 17 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR 52 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM 28 delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER 91 employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems significantly enhance clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
Rare Disease Identification and Management
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns 215. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. 16 developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO 72 utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. 216 created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. Some authors 217 have also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can effectively bridge knowledge gaps in rare disease management, connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
Clinical Report Generation
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care 218. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation 219. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption 220.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs 7. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 221, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors 222 proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines 223 have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG 224 employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors 225 have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
Precision Medicine Applications
Genomic Medicine
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity 226.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT 87 utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. 18 developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO 72 connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM 81 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through sophisticated retrieval mechanisms, these RAG systems significantly improve the clinical application of genetic information, such as disease inference 18 and phenotypic and genotypic analysis 72, while supporting more personalized treatment decisions in the emerging era of genomic medicine.
Personalized Health Management
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status 227. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE 29 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE 85 implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System 50 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP 92 utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
Healthcare Education
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences 228. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT 89 leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT 153 transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. 186 developed targeted applications in ophthalmology , including "EyeTeacher" for generating educational questions and "EyeAssistant" for clinical support, both enhanced through the integration of clinical guidelines and peer-reviewed documents for retrieval. In nursing education, Zhao et al 229 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through sophisticated retrieval mechanisms, these RAG-based education systems serve as valuable educational aid tools for healthcare practice skill training.
Clinical Research
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. 230.
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT 34 implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt 62 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and ultimately accelerating the translation of research findings into clinical practice.
Challenges and Future Direction
Trustworthiness
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans 1, 231. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source 181. Although certain re-ranking methodologies described in Section Reranker can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications 139. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora 2, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
Multilinguality
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored 232, 233. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages 234, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
Multimodality
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray 203, 235, 236), time series (e.g. ECG 152), sound recordings (e.g. lung sounds recorded 237), and video 238. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes 239. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the Biomedical domain and improving its accuracy and trustworthiness in the future.
Restricted Computing Resources
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge 1. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs 240.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks 241, 242. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts 243. Models such as QWQ 160 and Mistral Small 3.1 40 demonstrate competitive performance despite having significantly reduced parameter counts. Applying Portable LLMs 244 as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
Privacy
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential 245. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation 246 and the Health Insurance Portability and Accountability Act 247. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.
Conclusion
In conclusion, this comprehensive survey has provided a detailed overview of the application of RAG in the biomedical domain. It has examined the basic framework of biomedical RAG and discussed its retriever, reranker, and generation modules, as well as explored KG-enhanced biomedical RAG and multi-modal biomedical RAG. The survey has also explored clinical applications where biomedical RAG can be applied and reviewed recent studies in this field. Additionally, the paper has identified key challenges, including trustworthiness, multilingual capabilities, multimodal integration, limited computational resources, and privacy concerns. Overall, this work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems, ultimately contributing to the development of more accurate, accessible, and impactful healthcare AI systems that can meaningfully enhance patient care through evidence-based information access and support.