### v3

# Retrieval-Augmented Generation in Biomedicine: A Survey of Technologies, Datasets, and Clinical Applications

Abstract
Recent advances in large language models (LLMs) have demonstrated remarkable capabilities in natural language processing tasks. However, their application in the biomedical domain presents unique challenges, particularly regarding factual accuracy and up-to-date knowledge integration. Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these challenges by combining the generative capabilities of LLMs with external knowledge retrieval. This comprehensive survey examines the application of RAG in the biomedical domain, focusing on its technological components, available datasets, and clinical applications. We present a systematic analysis of retrieval methods, ranking strategies, and generation models, while also exploring the challenges and future directions in this rapidly evolving field. Our work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems and identifies key areas for future research and development.
Keywords
Biomedical RAG, Large Language Model, Retrieval-Augmented Generation, Information Retrieval, Natural Language Processing, Biomedical NLP
Introduction
Recent years have witnessed remarkable advancements in large language models (LLMs), resulting in a profound impact on various natural language processing (NLP) tasks Fan2024RAG, he2023survey. These models have demonstrated unprecedented capabilities in generating and understanding human-like text, achieving outstanding performance across NLP tasks, such as summarization, question answering, and information retrieval. The exceptional performance of LLMs in core NLP tasks has catalyzed their exploration in the biomedicine domain, where they show promise in supporting clinical decision-making, enhancing patient care quality, and improving clinical outcomes Wang2024JMLR.
However, the application of LLMs in the biomedical domain faces two critical challenges. First, these models often generate plausible-sounding but factually incorrect responses, a phenomenon commonly known as hallucination li2025biomedrag. This is particularly concerning in medical applications, where accuracy is paramount for patient safety. Second, the static nature of LLM training means that once the training process is complete, the model parameters remain fixed, resulting in an inability to access or incorporate up-to-date medical knowledge XIAO2025102888. This limitation is especially problematic in the rapidly evolving field of medicine, where new research findings and treatment guidelines emerge regularly.
Biomedicine RAG Dataset and Evaluation § IV Technology § III Preliminary § II Introduction § I Applications § V Discussion § VI Conclusion § VII Reranker Retriever Generation Domain Adaptation Medical Task Datasets Knowledge Source Datasets Evaluation Metrics Precision Medicine Applications Clinical Report Generation Clinical Decision Support Systems Healthcare Education Clinical Research Summary of dataset and evaluation approaches Summary of technological approaches Summary of RAG applications in biomedicine Challenges and Future Direction Knowledge Graph Enhanced Retrieval Dense Retrieval Sparse Retrieval Multi-modal Retrieval Commercial LLMs Open-Source LLMs Biomedical Specialized LLMs Retriever Stage Adaptation Reranker Stage Adaptation LLM Stage Adaptation Diagnostic and Treatment Decision Support Medical Question Answering Rare Disease Identification and Management Genomic Medicine Personalized Health Management
Figure 1: Structure of this paper
Retrieval Augmented Generation (RAG) has emerged as a promising solution to address these critical challenges frisoni2024generate. By combining the generative capabilities of LLMs with dynamic access to external knowledge from trusted sources, RAG systems can: (1) provide transparent rationales for generated responses by explicitly referencing source documents; (2) maintain access to current biomedical knowledge through real-time information retrieval; (3) reduce hallucinations by grounding responses in verified medical literature, thereby enhancing the accuracy and reliability of medical information processingFan2024RAG, he2023survey, XIAO2025102888.
The implementation of RAG in the biomedical domain presents unique challenges and opportunities. Biomedical information is characterized by complex terminology, specialized knowledge structures, and strict requirements for accuracy and reliability xiong2024benchmarking. Furthermore, the integration of RAG systems into clinical workflows requires careful consideration of factors such as information privacy, regulatory compliance, and clinical validation 10.1007/978-981-97-9437-9_2, santra2024regulatory. Therefore, a systematic survey of existing studies applying RAG in the biomedicine and healthcare domain is necessary to understand how these challenges are being addressed and to identify current research trends and gaps.
We considered papers that are directly related to the application of RAG techniques in the biomedicine or healthcare domain. Eligible studies include those that incorporate RAG components—such as retrievers, rerankers(optional), and generation models—or apply RAG to improve model performance on biomedical tasks. Specifically, we included works applying RAG to biomedical named entity recognition (NER), biomedical question answering (QA), multiple-choice reasoning, diagnostic classification, and medical dialogue systems, including, but not limited to, these areas. Studies were required to demonstrate how RAG integration enhances the performance, accuracy, reliability, or factual grounding of biomedical models, systems, or applications. Studies that use RAG only as a baseline or comparative model without substantively integrating RAG into their methodological framework were excluded. To collect relevant literature, we systematically searched PubMed and Google Scholar for studies published between 2022 and 2025, ensuring coverage of the most recent developments in biomedical RAG research. In total, 143 papers met the inclusion criteria and were included in this survey.
Concurrent to our work, several surveys have examined either biomedical LLMs or general-domain RAG systems. Surveys such as Liu et al. liu2024survey and He et al. he2023survey focus primarily on medical LLMs, emphasizing model evolution, training data, clinical applications, and safety considerations, but they do not analyze the RAG pipeline or retrieval–generation integration. Xiao et al. XIAO2025102888 further expand to multimodal LLMs in medicine, yet similarly concentrate on model architectures and clinical use cases rather than grounding mechanisms or retrieval-oriented methodologies.
In contrast, recent surveys on RAG—such as Fan et al. Fan2024RAG and Gao et al. Gao2023RAGSurvey—provide systematic overviews of RAG architectures, training paradigms, and evaluation practices in general domains. However, they do not address the unique challenges of biomedical knowledge sources, medical-language characteristics, domain adaptation strategies, or multimodal medical data.
Our survey bridges this gap by offering the first retrieval-centric and pipeline-level review of RAG techniques specifically tailored for biomedical applications. We synthesize advances from sparse/dense/LLM-based retrieval, biomedical rerankers, and domain-adapted generation models, and further integrate developments in knowledge graphs and multimodal retrieval. By unifying these threads, our work positions biomedical RAG as a distinct research area and provides a structured roadmap that helps developers and researchers understand current progress and effectively contribute to future advances. Table 1 provides a brief comparison of recent related survey.
Table 1: Comparative summary of related survey papers
Reference Year RAG Biomedical Pipeline Coverage Key Contributions
He et al. he2023survey 2023 Generation only Provides an overview of medical LLM development, applications, and ethical considerations, but does not address retrieval or grounding mechanisms.
Liu et al. liu2024survey 2024 Generation only Covers Med-LLM technologies, applications, and trustworthiness, yet lacks discussion of retrieval augmentation or biomedical evidence integration.
Xiao et al. XIAO2025102888 2025 Generation only Reviews LLMs and MLLMs in medicine, emphasizing model architecture and clinical scenarios, but does not examine retrieval, reranking, or grounding.
Fan et al. Fan2024RAG 2024 Retrieval + Generation Provides a structured overview of RAG architectures and training strategies in general domains, without biomedical datasets or domain-specific methods.
Gao et al. Gao2023RAGSurvey 2023 Retrieval + Generation Introduces Naive, Advanced, and Modular RAG frameworks, but does not cover biomedical tasks, knowledge sources, or clinical evaluation.
Ours 2025 Full pipeline: retrieval, reranking, generation, domain adaptation Offers the first domain-specific, pipeline-level survey of biomedical RAG, integrating sparse/dense/LLM-based retrievers, biomedical rerankers, domain-adapted generators, knowledge-graph RAG, multimodal retrieval, and medically grounded evaluation practices.
Overall, this survey provides the following key contributions:
  1. 1.
    A comprehensive review of retrieval-augmented technologies tailored to the biomedical domain, covering sparse, dense, LLM-based, and domain-specialized retrievers, as well as biomedical rerankers and generation models.
  2. 2.
    A detailed examination of medical datasets and evaluation, considering both knowledge source datasets and medical task benchmarks, along with their evaluation methods.
  3. 3.
    A practical, pipeline-oriented framework that examines existing biomedical RAG applications and offers developers and practitioners a clear roadmap for system implementation, while also highlighting future challenges and research directions.
Fig. 1 illustrates the overall structure of this survey. The remainder of the paper is organized as follows. Section II introduces the background and fundamental framework of biomedical RAG systems. Section III details the core technical components, including retrieval methods, ranking algorithms, and generation models. Section IV reviews the available datasets and evaluation methodologies. Section V discusses biomedical applications that leverage RAG. Section VI summarizes the key observations and discusses current challenges and future directions. Finally, Section VII concludes the paper.
Preliminary
To provide a clear conceptual and structural understanding of biomedical RAG systems, we first summarize their general workflow and key components. As Fig. 2 shows, in a typical biomedical RAG workflow, the system can be formally characterized as a composition of sequential functions operating over a query and an external knowledge source. Given a user query $q$ and a heterogeneous knowledge collection $K$, the end-to-end RAG process generates a response $y$ through the functional pipeline:
$y=G\big(q,\,S\big(q,\,R(q,K)\big)\big),$ (1)
where each component $R$, $S$ and $G$ represents a distinct stage of the retrieval–generation paradigm.
First, the Retriever $R(q,K)$ identifies a set of candidate documents relevant to the query and returns a document collection $D=\{d_{1},d_{2},\ldots,d_{n}\}\subseteq K$. Next, the Re-ranking Method $S(q,D)$ is an optional component of the RAG pipeline that refines the retrieved set by sorting and prioritizing documents according to task-specific relevance criteria, producing a re-ranked subset $D^{\prime}=[d^{\prime}_{1},d^{\prime}_{2},\ldots,d^{\prime}_{m}]$ with $D^{\prime}\subseteq D$. The use of this component, as well as the choice of scoring strategy, varies across biomedical applications. Finally, the Generation Function $G(q,D^{\prime})$ synthesizes the retrieved evidence with the original query to produce the final output:
$y=G(q,D^{\prime})=\text{LLM}\big(\text{prompt}(q,D^{\prime})\big),$ (2)
where the prompting function integrates the query with the selected context, $\text{prompt}(q,D^{\prime})=q\oplus D^{\prime}$. This formulation describes RAG as a modular and extensible computational paradigm, whose behavior can be adapted to a wide range of biomedical tasks through different designs of the generation function $G$.
To illustrate this flexibility, we formalize 4 representative instantiations of $G$ in biomedical applications:
  1. 1.
    Biomedical Text Classification. For classification-oriented tasks, the generation function $G$ selects the most probable label from the label set $\mathcal{C}$ under the model’s conditional distribution:
    $G_{\text{cls}}(q,D^{\prime})=\arg\max_{c\in\mathcal{C}}\,p_{\theta}(c\mid\text{prompt}(q,D^{\prime})).$ (3)
    This formulation encompasses a variety of biomedical classification scenarios, including diagnostic classification and multiple-choice question answering.
  2. 2.
    Biomedical Question Answering. For free-text biomedical QA, $G$ generates an answer sequence $y=[y_{1},\ldots,y_{T}]$ by sampling (or decoding) tokens from the conditional language model distribution:
    $G_{\text{qa}}(q,D^{\prime})=\arg\max_{y}\,p_{\theta}(y\mid\text{prompt}(q,D^{\prime})),$ (4)
    where $y$ represents the generated natural-language answer.
  3. 3.
    Clinical Dialogue or Consultation Systems. For multi-turn dialogue, $G$ conditions on the dialogue history $h$ in addition to the retrieved evidence:
    $G_{\text{dialog}}(q,D^{\prime},h)=\arg\max_{y}\,p_{\theta}(y\mid h\oplus\text{prompt}(q,D^{\prime})),$ (5)
    where $y$ is the next system response.
  4. 4.
    Biomedical Named Entity Recognition (NER). For sequence labeling tasks, $G$ predicts a sequence of token-level labels:
    $G_{\text{ner}}(q,D^{\prime})=\big[\arg\max_{y_{t}\in\mathcal{L}}\,p_{\theta}(y_{t}\mid\text{prompt}(q,D^{\prime}),t)\big]_{t=1}^{T},$ (6)
    where $\mathcal{L}$ is the label set and $T$ is the input sequence length.
These formulations highlight how biomedical RAG systems form a unified retrieval–generation pipeline, while tailoring the generation function to meet the requirements of diverse downstream tasks.
Beyond the task-specific instantiations of the generation function $G$, the RAG paradigm further benefits from the inherent extensibility of its external knowledge source. The knowledge repository $K$ is designed to be continuously updatable, enabling biomedical RAG systems to maintain timely access to newly emerging medical evidence. As new documents or knowledge collections—such as updated clinical guidelines, recently published biomedical literature, or domain-specific databases—become available, they can be seamlessly incorporated into the existing repository. Formally, the expanded knowledge source is expressed as $K_{\text{updated}}=K\cup K_{\text{new}}$, where $K_{\text{new}}$ denotes the newly added knowledge. This extensibility ensures that retrieval and generation remain grounded in the most current biomedical information, thereby enhancing the robustness and adaptability of RAG across diverse clinical and research applications.
Figure 2: Basic biomedical RAG framework.
RAG has been widely applied across diverse biomedical and healthcare scenarios, including clinical decision support, healthcare education, clinical report generation, and other related applications. Across these applications, RAG use diverse data sources-including health records johnson2016mimic, herrett2015cprd, scientific literature specter2020cohan, clinical guidelines Bashir2024, and genomic databases Lu2025BoostingGPT—to generate evidence-based responses that significantly improve both clinical care and research outcomes. A more detailed discussion of specific applications is provided in Section V.
Technology
In this section, we explore the technological underpinnings of biomedical RAG systems, which integrate specialized components for retrieval, re-ranking, and generation to effectively process domain-specific information. Our examination of retrieval mechanisms encompasses a range of approaches, including sparse Jin2022, zhang2024online and dense retrieval methods parmanto2024calm, Guo2024, LLM-based retrieval li2025biomedrag, ong2024surgeryllm, knowledge graph enhancement techniques Kakalou2024, Ascle, and emerging multi-modal retrieval capabilities designed to handle the diverse data formats prevalent in biomedical contexts Sun2024FactAware, Raminedi2024MultimodalTransformer, Upadhyaya2024. We further analyze re-ranking methodologies that refine initial retrieval results to prioritize the most relevant documents jin2023medcpt, jia2024podgpt, help address the challenge of relevance and correctness in biomedical information retrieval, where accuracy directly impacts clinical decision-making and research outcomes. Finally, we discuss the generation component of biomedical RAG systems, comparing closed-source OpenAI2025TextGeneration, Anthropic2025GettingStarted, GoogleGemini2025 and open-source models MetaAI2024LLaMA3, MosaicML2023Introducing, mistral_small_24b_2024, deepseekr1_2024 alongside domain adaptation techniques frisoni2024generate, llm-regulatory-framework that enable these general-purpose language models to produce specialized biomedical content with appropriate terminology, and contextual understanding.
Retriever
Retrievers are one of the core components of any RAG system, functioning as the mechanism that identifies and extracts relevant information from vast knowledge repositories in response to specific queries Gao2023RAGSurvey. In the biomedical domain, retrievers face distinctive challenges due to the heterogeneous nature of data structures—ranging from densely technical academic papers and specialized textbooks to complex biomedical knowledge graphs containing intricate entity relationships 10.1145/3648471, 10.1145/3462476. The domain’s specialized terminology, hierarchical concepts, and the life-critical nature of biomedical information make it more complex than information retrieval in general domainssivarajkumar2024clinical. Effective biomedical retrievers must navigate these complexities while maintaining high precision to support accurate downstream analysis and generation. In this subsection, we present four main retrieval paradigms used in biomedical RAG systems. Sparse and dense retrieval represent the most commonly adopted approaches, whereas knowledge graph–enhanced and multimodal retrieval are designed to handle more complex data structures, such as structured knowledge graphs and multimodal biomedical information.
Sparse Retrieval
Sparse retrieval techniques represent a foundational approach in information retrieval, particularly valuable in biomedical RAG systems which rely on terminology precision and exact keyword matching. These methods, exemplified by algorithms like BM25 (Best Matching 25) bm25 and TF-IDF (Term Frequency-Inverse Document Frequency) Jones1972, operate on the principle of lexical matching, where documents are indexed based on term occurrences rather than semantic meaning. Despite the rise of dense retrieval methods, sparse retrievers maintain their relevance due to their interpretability, computational efficiency, and effectiveness, when dealing with highly specialized vocabulary that characterizes biomedical literature.
Table 2 shows some recent publications that utilize sparse retrieval methods in biomedical RAG systems. These papers demonstrate various applications of sparse retrieval, from pure BM25 Das2025, Gong2024 implementations to hybrid approaches Lee2024, ALONSO2024102938, Abdullahi2024, Kim2024 that combine sparse and dense retrieval techniques. Notably, sparse retrieval is frequently treated as either a baseline for comparison or strategically incorporated as a component in ensemble solutions to increase retrieval diversity but also to improve generalization kamalloo:2024. Many of these works show that despite the attention given to newer neural approaches, traditional sparse retrieval methods remain competitive and often complementary to dense retrievers, particularly for specialized medical terminology and exact matching requirements xu2024bmretriever, Hu2024, izacard2021contriever. Their continued relevance in state-of-the-art biomedical RAG systems highlights the value of integrating lexical matching with semantic understanding when dealing with the precise terminology and complex nomenclature characteristic of biomedical literature.
In this line of work, BM25 stands out as the most widely adopted algorithm, featuring in more than half the studies (n=9) either independently or within hybrid systems. Established search tools including ElasticSearch ElasticsearchDocs, Whoosh WhooshDocs, and Entrez API EntrezAPI also prove popular, offering accessible implementation options without requiring extensive customization. A notable trend is the increasing development of hybrid approaches that combine BM25 with dense retrieval techniques xiong-etal-2024-benchmarking, Lee2024, ALONSO2024102938, Abdullahi2024, Kim2024, leveraging both lexical precision and semantic understanding. The consistent appearance of sparse retrieval methods across all publication years, including the most recent 2025 studies, demonstrates their enduring value alongside newer neural approaches in cutting-edge biomedical information retrieval systems.
Table 2: Sparse retrieval methods in biomedical RAG systems
Method Name Retrieval Method Task Year
Clinfo.ai Lozano2024ClinfoAI Entrez API Medical QA 2023
MEDRAG toolkit xiong-etal-2024-benchmarking Hybrid (BM25, SPECTER, Contriever, MedCPT) Medical QA 2024
SeRTS Hu2024 Self-Rewarding Tree Search based BM25 Biomedical QA (BioASQ-QA) 2024
Dual RAG System Lee2024 Hybrid (Dense + BM25 with language-specific tokenizers) Diabetes management 2024
MedExpQA ALONSO2024102938 BM25 and MedCPT Medical Question Answering 2024
CliniqIR Abdullahi2024 BM25 and MedCPT Diagnostic decision support 2024
VAIV Bio-Discovery Kim2024 Hybrid (Neural search + BM25) Biomedical knowledge discovery 2024
Genomics analysis GPT Lu2025BoostingGPT Azure AI Search with keyword matching Genetic variant annotation 2025
Customized GPT Gong2024 BM25 algorithm Medical QA 2025
Two-Layer RAG Das2025 Whoosh (Okapi BM25F ranking) Medical question answering 2025
LITURAt Peasley2025 Entrez API for PubMed retrieval Scientific data analysis 2025
Dense Retrieval
Dense retrievers represent a powerful approach to information retrieval that encodes queries and documents into continuous vector spaces to enable semantic matching based on learned representations rather than lexical overlap 10.1145/3637870. These retrievers have become important in the biomedical domain where terminology is complex and specialized, often requiring a nuanced understanding of context and meaning beyond simple keyword matching sivarajkumar2024clinical. For instance, synonyms like ‘cancer’ and ‘neoplasm’ refer to the same concept, while hierarchical relationships exist between terms such as ’infectious disease’ and its specific type ’COVID-19’, further complicating accurate information retrieval. To analyze biomedical RAG systems, we categorize dense retrievers into five distinct types: Type 1: Pre-trained language models (PLMs); Type 2: Commercial embedding APIs; Type 3: Domain-adapted PLMs; Type 4: LLM-based embeddings; and Type 5: Domain-specific retrievers. In the following paragraphs, we provide a detailed analysis of each retriever type’s characteristics, followed by a comprehensive review of recent works that implement these approaches in biomedical RAG systems.
Type 1: PLMs. PLMs are trained on general corpora and are widely used in various information retrieval systems to create dense representations of queries and documents, so-called embeddings. There are numerous open-source general embedding models available, which effectively capture semantic relationships and enable meaningful text matching, making them common foundational components in biomedical RAG systems. As shown in Table LABEL:tab:Classification, transformer-based PLMs like BERT, SBERT santander2022semantic, and Sentence Transformer thakur2020augmented have been widely adopted in recent biomedical retrieval systems. The continued popularity of these models is evident from numerous recent implementations. Additionally, newer embedding models such as BGE-embedding bge_embedding, Dense Passage Retriever Karpukhin2020 and GIST Large Embedding solatorio2024gistembed have gained significant adoption in recent literature, further diversifying the range of options available for biomedical retrieval tasks. Several other widely-used PLMs, including FastText bojanowski2017enriching, RoBERTa liu2019roberta, and SimCSE gao-etal-2021-simcse, though not specifically represented in Table LABEL:tab:Classification, remain viable alternatives for developing effective biomedical RAG applications.
Table 3: Classification of dense retriever types in biomedical RAG systems
Method Retriever Task Year
Type 1: PLMs
AskFDALabel llm-regulatory-framework Sentence Transformer thakur2020augmented FDA drug labeling extraction 2024
CaLM parmanto2024calm BGE-embedding bge_embedding with Chroma DB Supporting caregivers FM 2024
RAG-HPO garcia2024deepphenotyping FastEmbed fastembed Rare genetic disorders automated deep phenotyping 2024
RALL Guo2024 Dense Passage Retriever Karpukhin2020 Lay language generation 2024
RT li2024rt BERT embeddings devlin2019bert Few-shot medical NER 2024
RAG-GPT xu2024evaluation BGE-embedding bge_embedding Breast cancer nursing care QA 2024
GNQA darnell2024biomedical HNSW graphs malkov2016efficient Medical QA 2024
Klang et al. klang2024assessing GIST Large Embedding solatorio2024gistembed with FAISS ICD-10-CM coding 2024
RAGPR zheng2025integrating SBERT embeddings santander2022semantic Personalized physician recommendations 2025
JGCLLM Fukushima2025 GLuCoSE-base-ja vectors Fukuchi2023 Genetic counseling support 2025
Type 2: Commercial APIs
DRAGON-AI toro2024dragonai OpenAI Text-Embedding openai_embeddings with Chroma DB Ontology generation 2024
RISE wang2024diabetesrag OpenAI Text-Embedding with FAISS Diabetes-related inquiries 2024
Dual RAG lee2024dualrag UPstage API and OpenAI Text-Embedding Diabetes management 2024
FAVOR-GPT li2024favorgpt Open AI Text-Embedding with Weaviate DB Genome variant annotations 2024
Endo-chat fu2024application Open AI Text-Embedding with Faiss medical QA for gastrointestinal endoscopy 2024
ChatENT long2024chatent Open AI Text-Embedding Medical QA in otolaryngology 2024
GuideGPT steybe2025evaluation Open AI Text-Embedding MRONJ QA on prevention, diagnosis, and treatment 2024
RECTIFIER unlu2024retrieval Open AI Text-Embedding with Faiss Clinical trial screening for heart failure patients 2024
RAP azimi2025evaluation Amazon Titan Text Embeddings v2 aws_titan_embedding Nutrition-related question answering 2025
InfectA-Chat selcuk2025infecta Open AI Text-Embedding Infectious disease monitoring and QA 2025
DCRAG kainer2025effectiveness OpenAI Text-Embedding Auto-annotation of plant phenotype 2025
Type 3: Domain adaptation PLMs
PM-Search Jin2022 BioBERT 10.1093/bioinformatics/btz682 Clinical literature retrieval 2022
LADER LADER PubMedBERT pubmedbert Biomedical literature retrieval 2023
CLEAR lopez2025clinical BioBERT Clinical information extraction 2025
KG-RAG Soman2024 PubMedBERT Biomedical Multiple-choice questions 2024
WeiseEule Aftab2024 MedCPT, BioBERT, BioGPT luo2022biogpt with Pinecone DB Biomedical QA 2024
Type 4: LLM-based embedding
BiomedRAG li2025biomedrag MedLLaMA 13b Biomedical NLP tasks 2024
SurgeryLLM ong2024surgeryllm GPT4All anand2023gpt4all with Chroma DB Surgical decision support 2024
Myers et al. Myers2025 LLM2Vec (for comparison) llm2vec Clinical Information retrieval 2024
Type 5: Domain adaptation retriever
Self-BioRAG jeong2024improving MedCPT jin2023medcpt Medical QA 2024
WeiseEule Aftab2024 MedCPT, BioBERT, BioGPT with Pinecone DB Biomedical QA 2024
CliniqIR Abdullahi2024 MedCPT and BM25 Diagnostic decision support 2024
MEDRAG Aftab2024 BM25, SPECTER specter2020cohan, Contriever izacard2021contrieverand MedCPT Biomedical RAG Tool 2024
RAMIE zhan2025ramie MedCPT, Contriever and BMRetriever xu2024bmretriever Biomedical IR about dietary supplements 2025
Type 2: Commercial embedding APIs. The proliferation of foundation models has led to the emergence of proprietary embedding solutions offered by major AI private organizations. These commercial embedding APIs have gained significant adoption in biomedical research, despite their associated costs, due to several compelling advantages: i) they typically leverage substantially larger pretraining datasets, ii) benefit from extensive computational resources during development, and iii) receive continuous refinements based on diverse user interactions. As shown in Table LABEL:tab:Classification, OpenAI’s embedding models dominate this category in recent studies, underscoring their enhanced semantic representation capabilities. While OpenAI’s offerings currently dominate, alternative solutions are also emerging. For example, the RAP system azimi2025evaluation utilizes Amazon’s Titan Text Embeddings v2 aws_titan_embedding for nutrition-related question answering, and other commercial APIs, such as Anthropic’s Voyage-2 anthropic_embeddings and Google’s Vertex AI embeddings google_vertexai_embeddings, though not represented in Table LABEL:tab:Classification, can also be employed in biomedical RAG applications
Type 3: Domain adaptation PLMs. Domain-adapted PLMs are language models that have been specifically pre-trained or fine-tuned on biomedical and clinical corpora, allowing them to develop specialized representations of medical terminology, concepts, and relationships 10.1093/bioinformatics/btz682, pubmedbert. These models mitigate a fundamental challenge in biomedical information retrieval: general-domain language models often struggle with the highly specialized vocabulary, complex semantic relationships, and unique linguistic patterns prevalent in biomedical literature Lyu2024LanguageMA. By adapting pre-training to include PubMed articles, clinical notes, or other domain-specific corpora, these models develop a more nuanced understanding of biomedical language, resulting in competitively improved performance on retrieval tasks involving medical concepts and relationships compared to their general-domain counterparts Lyu2024LanguageMA. As shown in Table LABEL:tab:Classification, examples of these domain-adapted models include PubMedBERT pubmedbert, BioBERT 10.1093/bioinformatics/btz682, and BioGPT luo2022biogpt, which have been widely adopted in biomedical RAG systems. For instance, PM-Search Jin2022 utilized BioBERT for clinical literature retrieval, effectively leveraging the model’s domain-specific representations to improve search precision. Similarly, PubMedBERT has been successfully deployed in LADER LADER for biomedical literature retrieval and in KG-RAG Soman2024 for answering biomedical multiple-choice questions, demonstrating its versatility across different tasks. More recent systems like WeiseEule Aftab2024 compare multiple biomedical models in their study, including BioBERT, BioGPT, and MedCPT jin2023medcpt, to achieve enhanced performance in biomedical QA.
Type 4: LLM-based embedding. Generative LLMs have demonstrated remarkable generative capabilities across diverse NLP tasks, owing to their sophisticated understanding of language and context. With billions of parameters and training on vast corpora, these models develop rich internal representations that capture nuanced semantic relationships and contextual dependencies in text Zhao2023LLMSurvey. Recognizing this powerful language comprehension ability, researchers have begun leveraging open-source LLMs as embedding generators for retrieval tasks, extracting representations from their internal layers to create semantically meaningful document and query embeddings. As shown in Table LABEL:tab:Classification, notable implementations include BiomedRAG li2025biomedrag, which utilizes MedLLaMA-13B wu2023pmc; SurgeryLLM ong2024surgeryllm, which employs GPT4All anand2023gpt4all; and the comparative study by Myers et al. Myers2025, which evaluates LLM2Vec llm2vec for clinical information retrieval applications. However, most LLMs are decoder-only architectures that employ causal attention mechanisms, which inherently limit their ability to generate rich contextualized representations. Thus, researchers are increasingly focusing on developing more advanced LLM-based embedding methods, such as Llama2Vec li-etal-2024-llama2vec, Landmark Embedding luo-etal-2024-landmark, which provide promising dense retrieval alternatives in biomedical RAG.
Type 5: Domain adaptation retrievers. Although the aforementioned off-the-shelf retrievers are readily available, the task of searching for relevant and accurate documents in the medical domain remains challenging. Consequently, customized retrievers tailored specifically to scientific and biomedical tasks have been developed. As shown in Table LABEL:tab:Classification, these advanced retrievers include Contriever izacard2021contriever, MedCPT jin2023medcpt, SPECTER specter2020cohan, BMRetriever xu2024bmretriever, and the MEDRAG toolkit li2025biomedrag. These specialized retrieval architectures demonstrate the importance of domain adaptation in biomedical information retrieval, significantly improving retrieval performance in several studies jeong2024improving, Abdullahi2024, zhan2025ramie. For example, Jeong et al. jeong2024improving proposed Self-BioRAG, which leverages MedCPT to improve medical question answering through a self-augmentation approach. Similarly, the MEDRAG framework Aftab2024 combines multiple retrieval strategies, including BM25, SPECTER, Contriever, and MedCPT to create a comprehensive biomedical RAG tool. RAMIEzhan2025ramie utilizes MedCPT, Contriever, and BMRetriever as retrievers, builds a RAG framework to extract diverse types of information about dietary supplements (DSs) from clinical records, enhanced overall accuracy in its experiments zhan2025ramie. These implementations highlight how domain-adapted retrievers benefit biomedical RAG systems.
Table 4: Common vector stores in biomedical RAG
Store Description
Chroma chroma2023 AI-native open-source vector database with built-in functionality
Faiss faiss2023 Library for efficient similarity search and clustering of dense vectors
Pinecone pinecone_overview Production-ready vector database for similarity search at scale
Weaviate weaviate_developers Graph-based vector search engine with semantic capabilities
HNSW malkov2016efficient Hierarchical navigable small world graphs for approximate nearest neighbor search
Dense Retriever Infrastructure. Vector stores provide an infrastructure element for dense retrievers in semantic search systems han2023comprehensive. These specialized databases efficiently index, store, and search high-dimensional vector embeddings, offering optimized methods like K-nearest neighbors (KNN) for RAG developers. Vector databases enable rapid retrieval of semantically similar documents at scale, thereby improving the efficiency and reducing the response time of biomedical RAG systems. The process of embedding, storing, and searching documents involves transforming text into vector representations, organizing them in specialized data structures, and efficiently retrieving the most relevant matches. Based on our survey of biomedical RAG publications presented in Table LABEL:tab:Classification, we compiled the most commonly used vector databases in Table 4, among which FAISS and Chroma emerge as the most popular choices, being utilized 4 and 3 times, respectively.
Knowledge Graph Enhanced Retrieval
Knowledge graph retrievers provide structured, interlinked domain knowledge in biomedical RAG, helping to address several limitations of traditional retrieval methods. Unlike vector-based approaches, knowledge graphs capture explicit relationships between biomedical entities, enabling more precise retrieval of contextually relevant information, enhanced reasoning capabilities, and improved explainabilitypeng2024graph. Table 5 presents recent biomedical RAG studies that incorporate knowledge graphs, outlining their application domains and the knowledge graph sources they employ. The application areas covered by these systems include medical question answering Soman2024, knowledge extraction DiMaria2024, and diagnosis prediction Gao2025.
Table 5: Recent biomedical RAG studies based on knowledge graphs
Method Task Knowledge Graph Used Year
KG-RAG Soman2024 Biomedical text generation, medical question answering SPOKE biomedical knowledge graph 2024
HEALIE Kakalou2024 Personalized medical content generation HEALIE KG (customized) 2024
KRAGEN Matsumoto2024 Biomedical problem solving Alzheimer’s knowledge graph (AlzKB) 2024
Ascle Ascle Various medical text generation tasks UMLS 2024
DALK li2024dalk Alzheimer’s QA Alzheimer’s knowledge graph (AlzKB) 2024
Gilbert et al. Gilbert2024 Medical information structuring and interlinking Not specified 2024
NetMe 2.0 DiMaria2024 Biomedical knowledge extraction BKGs constructed using OntoTagMe and Wikidata 2024
NEKO XIAO202560 Knowledge mining in synthetic biology PubMed-derived knowledge graphs 2025
DR.KNOWS Gao2025 Diagnosis prediction from EHRs UMLS 2025
ESCARGOT Matsumoto2025 Biomedical reasoning and knowledge retrieval Alzheimer’s knowledge graph (AlzKB) 2025
Figure 3: The Framework of Knowledge Graphs-Based Biomedical RAG
To further clarify how knowledge graphs are incorporated into biomedical RAG systems, Fig. 3 illustrates the integration of knowledge graphs with LLMs through four key components:
Biomedical knowledge graph construction or curation involves the systematic process of building, updating, or adapting knowledge graphs from domain-specific ontologies, literature, and other data sources to create structured biomedical knowledge repositories. For example, NetMe 2.0 DiMaria2024 constructs biomedical knowledge graphs on-the-fly using OntoTagMe for entity linking, while NEKO XIAO202560 mines PubMed literature to build knowledge graphs for synthetic biology. By contrast, KG-RAG Soman2024 utilizes the pre-existing SPOKE biomedical knowledge graph.
Query structuration encompasses the conversion of natural language queries or user inputs into structured queries that efficiently retrieve relevant subgraphs from the knowledge graph. For instance, ESCARGOT Matsumoto2025 employs Cypher queries for precise retrieval, KRAGEN Matsumoto2024 vectorizes knowledge graphs for semantic searching, and DR.KNOWS Gao2025 transforms queries to retrieve UMLS-based diagnostic pathways aligned with patient data.
Path-based or entity-centric knowledge extraction refers to the process of extracting relevant knowledge patterns, paths, or entity-centered information from the knowledge graph based on the transformed query. Specifically, DR.KNOWS Gao2025 utilizes attention-based path ranking for diagnostic reasoning, KRAGEN Matsumoto2024 applies graph-of-thoughts for multi-hop relationships, and HEALIE Kakalou2024 extracts paths incorporating both medical knowledge and socio-economic factors.
Strategic integration into LLM prompts involves methods for incorporating retrieved knowledge into prompts in ways that maximize factual accuracy while optimizing token usage and computational efficiency. For example, KG-RAG Soman2024 implements token-optimized retrieval to reduce consumption, whereas ESCARGOT Matsumoto2025 selectively integrates knowledge through a dynamic Graph of Thoughts. In addition, NetMe 2.0 DiMaria2024 transforms graph data into natural language prompts via its Graph-RAG module.
Multi-modal Retrieval
Multimodal biomedical RAG represents an advanced approach to healthcare AI that integrates LLMs with multiple types of data inputs beyond text, such as medical images, electrocardiogram (ECG), and other clinical measurements. These systems can enhance diagnostic accuracy and clinical decision support by leveraging the complementary strengths of different data modalities, retrieving relevant information from specialized knowledge bases, and generating comprehensive insights that would be impossible to achieve with single-modality approaches abootorabi2025askmodalitycomprehensivesurvey. As shown in Table 6, multimodal biomedical RAG systems have been successfully deployed for amblyopia detection Upadhyaya2024, prescription label interpretation Thetbanthad2025, lung cancer staging Tozuka2024, digital pathology analysis Omar2024, and automated radiology report generation pmlr-v219-ranjit23a, Sun2024FactAware, xia2024mmedrag, Xia2024RULE, Yang2025STREAM.
Figure 4: Architectural workflow of multimodal biomedical RAG system
Table 6: Multimodal Biomedical RAG Systems
Model Task Modalities Year
Ranjit et al. pmlr-v219-ranjit23a Chest X-Ray report generation Chest X-ray and text 2023
FactMM-RAG Sun2024FactAware Radiology report generation X-Ray and text 2024
MMed-RAG xia2024mmedrag Medical VQA & Report generation Medical images and text 2024
MEAG Upadhyaya2024 Amblyopia diagnosis Eye tracking and text 2024
Tozuka et al. Tozuka2024 Lung cancer staging CT findings and TNM classifications 2024
GPT4DFCI-RAG Omar2024 Digital pathology QA Pathology images and text 2024
Raminedi et al. Raminedi2024MultimodalTransformer Radiology report generation X-ray images and text 2024
RULE Xia2024RULE Medical VQA & Report generation Medical images and text 2024
Thetbanthad et al. Thetbanthad2025 Prescription label identification Image (labels) and text 2025
Hu et al. Hu2025 Pathology report generation Whole slide images and text 2025
STREAM Yang2025STREAM Chest X-ray report generation Chest X-ray and text 2025
Figure 4 illustrates the architectural workflow of a multimodal biomedical RAG system. The key distinction between multimodal biomedical RAG and traditional biomedical RAG lies in data input diversity and processing architecture. While traditional RAG systems primarily process textual queries to retrieve text-based medical knowledge, multimodal RAG incorporates and processes non-textual data (images, sensor readings, etc.) alongside text, requiring specialized encoders for each modality to meaningfully integrate these diverse inputs abootorabi2025askmodalitycomprehensivesurvey. The basic framework of multimodal biomedical RAG typically consists of: (1) modality-specific encoders (e.g., Vision Transformers chen2021outperform for images, specialized encoders for time-series data chen2021outperform); (2) knowledge retrieval components that access domain-specific databases based on multimodal query representations; and (3) generation modules that synthesize comprehensive outputs informed by both the multimodal inputs and retrieved knowledge. For instance, Ranjit et al.pmlr-v219-ranjit23a, FactMM-RAG Sun2024FactAware and STREAM Yang2025STREAM use multimodal embeddings for retrieval of relevant candidate radiology text to generate Chest X-Ray report. MMed-RAG xia2024mmedrag and RULE Xia2024RULE build a multimodal RAG system for Medical Visual Question Answering (Medical VQA) and report generation. The MEAG Upadhyaya2024 system uses specialized encoders for eye movement data alongside text, Raminedi et al. Omar2024 employ Vision Transformers as encoders for radiological images with cross-attention mechanisms for fusion with text, while Thetbanthad et al. Thetbanthad2025 first converts prescription label images to text through OCR before retrieving relevant pharmaceutical information from DrugBank.
Reranker
Rerankers in biomedical RAG systems serve as intermediary components that refine and prioritize the relevance of initially retrieved documents before they are processed by the generative models sivarajkumar2024clinical. Positioned between retrievers and generators, rerankers apply appropriate ranking algorithms to re-score and re-order candidate documents based on their contextual relevance to the query, effectively functioning as a second-stage filtering mechanism that enhances the precision of information ultimately provided to the generation component. By prioritizing the most pertinent information, rerankers help to minimize hallucinations and improve the veracity of generated outputs.
Biomedical RAG rerankers can be categorized into distinct types based on their underlying methodologies. Statistical rerankers apply traditional information retrieval techniques; for instance, CliniqIR Abdullahi2024 and MEDRAG Aftab2024 both implement Reciprocal Rank Fusion (RRF) to combine multiple retrieval signals, while DRAGON-AI toro2024dragonai employs Maximal Marginal Relevance (MMR) to balance relevance with diversity. In addition, WeiseEule Aftab2024 introduces a keyword frequency–based ranking system tailored for biomedical literature retrieval. Model-based rerankers employ transformer architectures to perform contextual relevance assessment; specifically, PodGPT jia2024podgpt uses a general embedding model (bge-reranker-large) for scientific literature retrieval, whereas the CHA2DS2-VASc risk factor extraction system adejumo2024retrieval utilizes a specialized cross-encoder model known as BGE M3-Embedding chen2024m3embedding to enhance clinical document relevance scoring. Within this category, some domain adaptation retrievers incorporate a reranker as a component of their model; for example, MedCPT jin2023medcpt trains a specific reranker denoted as CrossEnc. Additionally, LLM-based rerankers leverage foundation models’ reasoning capabilities; GNQA darnell2024biomedical employs GPT-3.5/4/4o as zero-shot binary classifiers for relevance determination, and LmRaC craig2024lmrac implements an LLM-based paragraph usefulness assessment with a minimum threshold score of 7. (on a 10 point scale). Finally, hybrid approaches combine multiple strategies to enhance performance. For example, Clinfo.ai Lozano2024ClinfoAI combines LLM classification with BM25 to re-rank retrieved articles.
Generation
The generation component represents the last element in biomedical RAG systems, responsible for synthesizing retrieved information into human-like responses that address user queries. In biomedical contexts, this process requires balancing natural language fluency with domain-specific accuracy, as generated content often informs healthcare decisions lee2024enhancing. This section examines various generation models employed in recent biomedical RAG studies.
Current generation models employed in biomedical RAG systems can be categorized into three distinct classes based on their development approach and specialization: general-domain open-source models, commercial proprietary models, and biomedical-specialized models. These categories present different trade-offs in terms of accessibility, customization potential, and domain-specific performance. Table LABEL:tab:llm_taxonomy showcases representative models from each category and highlights recent studies that utilize these models. It is important to note that our analysis of biomedical-specialized LLMs focuses specifically on models trained directly on biomedical or healthcare corpora, deliberately excluding applications developed through prompt engineering or Chain-of-Thought (CoT) methodologies.
As shown in Table LABEL:tab:llm_taxonomy, the majority of LLMs employ decoder-only architectures, with T5 being a notable exception. This architectural preference stems from several inherent advantages for generative tasks, particularly the natural suitability for next-token prediction and significantly improved parameter efficiency Zhao2023LLMSurvey. Among open-source LLMs, LLaMa 2 HugoTouvron2023LLaMA2 is the most popular, with 15 biomedical applications built upon it. LLaMa 3 MetaAI2024LLaMA3, Mistral 7B mistral7b_2024, and T5 2020t5 are also widely used. These open-source LLMs have various sizes, from 0.06B (T5) to 671B (Deepseek R1), providing flexible options for building Biomedical RAG systems. Recently, several powerful general open-source LLMs have emerged, including, Mistral Small 3.1 mistral_small_24b_2024 and Gemma 3 gemma-3-27b-it. Although we have not found these newer models being used for biomedical RAG systems in our survey, we believe they show promise due to their outstanding performance in general domain tasks.
For commercial LLMs, the ChatGPT series APIs maintain a dominant position, presumably attributable to their superior performance. Nevertheless, several studies have demonstrated that Claude and Gemini achieve comparable efficacy when implemented in biomedical RAG applications wang2024diabetesrag, hewitt2024llmneuropathology, Lammert2024ExpertLLM, thereby offering viable alternatives to ChatGPT. Furthermore, it is noteworthy that both DeepSeek and Mistral have established dual accessibility paradigms: providing open-source parameters for their base models while simultaneously offering commercial API services.
In contrast to open-source and commercial LLMs, biomedical specialized LLMs have been relatively underutilized in biomedical RAG applications. Recent studies ozaki2024confidence, zhan2025ramie, alkaeed2025open have demonstrated that within RAG frameworks, biomedical specialized LLMs do not consistently outperform open-source alternatives with statistical significance. Thus, the more effective methodologies for leveraging biomedical specialized LLMs within retrieval-augmented generation frameworks remain an area requiring substantial further investigation.
Table 7: Taxonomic Classification of LLMs for Embedding Applications
Model Parameters Architecture Studies
Open-Source LLMs
T5 2020t5 0.06-11B Encoder-Decoder VAIV Kim2024, CLEAR lopez2025clinical, DR.KNOWS Gao2025
ChatGLM3 glm2024chatglm 6B Decoder-only COPD Copd-ChatGLM, ChatZOC luo2024retrieval
LLaMA 2 HugoTouvron2023LLaMA2 7-70B Decoder-only Guo et al. Guo2024, KREIMEYER et al. kreimeyer2024rag, CaLM parmanto2024calm, MMRAG zhan2025mmrag, Yu et al. pmlr-v225-yu23b, RAMIE zhan2025ramie, EyeGPT chen2024eyegpt, KG-RAG Soman2024, JMLR Wang2024JMLR, BiomedRAG li2025biomedrag, Self-BioRAG jeong2024improving, CaLM parmanto2024calm, Du et al. du2024cognitive, SurgeryLLM ong2024surgeryllm, Kreimeyer et al. kreimeyer2024rag
LLaMA 3 MetaAI2024LLaMA3 8-70B Decoder-only MMRAG zhan2025mmrag, Fatharanihttps et al. fatharani2025pharmacogenomics, RAMIE zhan2025ramie, Woo et al. woo2025custom, RAG-HPO garcia2024deepphenotyping, i-MedRAG xiong2025iterative, Hewitt et al. hewitt2024llmneuropathology, PodGPT jia2024podgpt, SurgeryLLM ong2024surgeryllm
Phi-3 Mini phi3mini4k_2024 3.8B Decoder-only Fatharanihttps et al. fatharani2025pharmacogenomics
Mistral 7B mistral7b_2024 7B Decoder-only Boulos et al. kamel2024chat, RAMIE zhan2025ramie, Woo et al. woo2025custom, RAGPR zheng2025integrating, LITURAt Peasley2025, Kreimeyer et al. kreimeyer2024rag
QWen qwen32b_2024 32B Decoder-only NEKO XIAO202560, Thetbanthad Thetbanthad2025
Baichuan wang2025baichuan 7-13B Decoder-only ChatZOC luo2024retrieval
Falcon almazrouei2023falconseries 7-180B Decoder-only AskFDALabel llm-regulatory-framework, CaLM parmanto2024calm
Zephyr tunstall2023zephyr 7B Decoder-only Moser et al. moser2024synthetic
AceGPT huang2024acegpt 7-13B Decoder-only InfectA-Chat selcuk2025infecta
Gemma team2024gemma 1-27B Decoder-only PodGPT jia2024podgpt
Deepseek R1 deepseekr1_2024 671B Decoder-only Feng et al. feng2025retrieval
Commercial LLMs
ChatGPT-3.5/4/4o OpenAI2025TextGeneration Proprietary Decoder-only Woo et al. woo2025custom, RAGPR zheng2025integrating, ChatZOC luo2024retrieval, KG-RAG Soman2024, Clinfo.ai Lozano2024ClinfoAI, Yu et al. pmlr-v225-yu23b, ChatENT long2024chatent, CaLM parmanto2024calm, Rau et al. rau2024retrieval, Endo-chat fu2024application, BiomedRAG li2025biomedrag, Lu Lu2025BoostingGPT, Tan an2025chatgpt, GNQA darnell2024biomedical, Puts puts2025icd10, DRAGON-AI toro2024dragonai, Choi Choi2024, RISE wang2024diabetesrag, Du et al. du2024cognitive,i-MedRAG xiong2025iterative , Hewitt et al. hewitt2024llmneuropathology, RECTIFIER unlu2024retrieval, Kainer et al. kainer2025effectiveness, Gong et al. Gong2024, VAIA Kim2024, RareDxGPT Zelin2024RareDisease, Selcuk et al. selcuk2025infecta, FAVOR-GPT li2024favorgpt, Markey et al. markey2025rags,
Claude-3.5 Anthropic2025GettingStarted Proprietary Decoder-only Woo et al. woo2025custom, RISE wang2024diabetesrag, Hewitt et al. hewitt2024llmneuropathology
Gemini GoogleGemini2025 Proprietary Multimodal Upadhyaya et al. Upadhyaya2024, Tozuka et al. Tozuka2024, MEREDITH Lammert2024ExpertLLM
Biomedical Specialized LLMs
MedAlpaca han2023medalpaca 7-13B LLaMA-based RAMIE zhan2025ramie
PMC-LLaMA Wu2024PMC-LLaMA 7B LLaMA-based RAMIE zhan2025ramie, BiomedRAG li2025biomedrag, Ozaki et al. ozaki2024confidence
BioMistral labrak2024biomistral 7B Mistral-based RAMIE zhan2025ramie, pRAGe buhnila2024retrieve
MEDITRON chen2023meditron70b 7-70B LLaMA-based Ozaki et al. ozaki2024confidence, Alkaeed et al. alkaeed2025open
Meerkat kim2024small 7B Mistral-based Sohn et al. sohn2024rationale
Domain Adaptation
Domain adaptation in biomedical RAG systems constitutes a methodological framework that can be strategically implemented across various RAG pipeline stages to optimize performance for medical applications siriwardhana2023improving. This approach facilitates enhanced performance on biomedical tasks while simultaneously preserving the fundamental capabilities inherent to foundation models. Based on our systematic analysis of the literature, we propose a taxonomy of domain adaptation techniques categorized according to their implementation within specific RAG pipeline components.
Retriever Stage Adaptation. Specialized retrieval components constitute the initial layer of domain adaptation, wherein techniques such as contrastive learning (implemented in Contriever izacard2021contriever and MedCPT jin2023medcpt) enable the development of more semantically rich biomedical embeddings. These adaptations augment the system’s capacity to identify and retrieve pertinent medical documents through enhanced comprehension of domain-specific terminology and conceptual relationships. Additional retriever adaptation methodologies include continued pre-training on biomedical corpora (as demonstrated in BMRETRIEVER xu2024bmretriever) and domain-specific architectural modifications (exemplified by LADER LADER with PubMedBERT initialization). These approaches substantially enhance retrieval precision for specialized medical literature, effectively mitigating the lexical and semantic disparities between general and biomedical language domains.
Reranker Stage Adaptation. The reranking phase presents a subsequent opportunity for domain specialization, employing methodologies such as the development of dedicated biomedical rerankers (e.g., MedCPT’s CrossEnc reranker jin2023medcpt) and the implementation of domain-specific scoring mechanisms (e.g., BiomedRAG’s chunk scorer li2025biomedrag). Notably, the aforementioned Model-based rerankers and LLM-based rerankers can be conceptualized as domain adaptation strategies operationalized within the reranking stage. These specialized rerankers, trained with domain-specific datasets and optimization objectives tailored to biomedical document ranking, improve the system’s ability to prioritize documents by capturing subtle domain-specific signals that general models often fail to detect, thereby enhancing retrieval precision for specialized medical queries.
LLM Stage Adaptation. At the generation stage, adaptation of the language model itself through methodologies such as fine-tuning (MEDGENIE frisoni2024generate, SeRTS Hu2024), parameter-efficient approaches like Low-Rank Adaptation (LoRA)hu2022lowrank (implemented in AskFDALabel llm-regulatory-framework), and continued pre-training (RALL Guo2024, ChatENT long2024chatent) enables more accurate synthesis and explication of biomedical information. JLMR Wang2024JMLR joint train LLM and retrieval model, which enhances biomedical RAG’s ability to retrieve clinical guidelines and leverage medical knowledge to reason and answer questions. These adaptations enhance the model’s comprehension of domain knowledge, facilitate the generation of appropriate medical terminology, and help improve factual accuracy in response generation.
This multi-level adaptation framework creates biomedical RAG systems that leverage the general capabilities of foundation models while incorporating the specialized knowledge representation and reasoning patterns required for healthcare and life sciences applications. Through the strategic application of adaptation techniques at each stage of the RAG pipeline, these systems achieve enhanced precision, improved contextual understanding, and more reliable generation of biomedical information, contributing to ongoing progress in biomedical information retrieval and synthesis.
Datasets and Evaluation
The evaluation and development of biomedical RAG systems rely heavily on appropriate datasets. In this section, we present a comprehensive overview of datasets commonly used in biomedical RAG research, categorizing them into two main types: knowledge source datasets and medical task benchmarks, along with their corresponding evaluation metrics. This categorization reflects the dual nature of RAG systems, which require both comprehensive knowledge bases and task-specific evaluation benchmarks.
Knowledge Source Datasets
The effectiveness of biomedical RAG systems considerably depends on the quality, comprehensiveness, and diversity of their knowledge sources gao2023retrieval, Fan2024RAG. These knowledge repositories serve as the informational backbone from which these systems retrieve and synthesize medical content. High-quality knowledge sources can improve the generated responses, making them not only relevant but also reflect the current biomedical consensus and best practices huang2024toolcalling, long2024chatent, jeong2024improving.
Biomedical knowledge spans multiple dimensions—from basic science research to clinical guidelines, from pharmaceutical data to standardized medical terminology. To address this complexity, effective biomedical RAG systems integrate diverse knowledge sources that complement each other in scope, specialization, and format. This integration enables systems to comprehensively address the multifaceted nature of medical queries Aftab2024, xu2024bmretriever.
Table LABEL:tab:med_knowledge shows the public knowledge source utilized in the surveyed studies. Among these sources, PubMed, UMLS bodenreider2004umls, PMC, and Medical textbooks jin2020disease demonstrate predominant adoption, primarily attributable to their comprehensive coverage of biomedical concepts and extensive document collections. Additionally, we notice that several studies Choi2024, du2024cognitive, zheng2025integrating have implemented biomedical RAG systems based on proprietary clinical data that is unpublished due to privacy concerns, while another subset of studies Lee2024, masanneck2025evaluating, xu2024evaluation, azimi2025evaluation, huang2024acegpt, sevgi2024medical focus on a specific medical specialty, wherein specialized medical guidelines about particular clinical specialties are leveraged for constructing a targeted knowledge source.
Table 8: Comprehensive Medical Knowledge Sources
Source Type Description Studies
MIMIC-IV johnson2023mimic EHR Decade of hospital admissions between 2008 and 2019 with data from about 300K patients and 430K admissions Moser et al. moser2024synthetic
PubMed Literature Repository of over 36 million biomedical research citations and abstracts, containing about 4.5B words Clinifo.ai Lozano2024ClinfoAI, MEDRAG xiong-etal-2024-benchmarking, Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, RISE wang2024diabetesrag, MEREDITH Lammert2024ExpertLLM, LITURAt Peasley2025, DR.KNOWS Gao2025, Aftab Aftab2024, CliniqIR Abdullahi2024, PM-Search Jin2022, Kreimeyer kreimeyer2024rag, Kim Kim2024
PMC Literature 8 million full-text biomedical and life sciences articles with free access, about 13.5B words Self-BioRAG jeong2024improving, Soong et al. soong2024improving, PodGPT jia2024podgpt
TripClick rekabsaz2021tripclick Literature Biomedical literature search and retrieval dataset LADER LADER
GNQA darnell2024biomedical Literature 3000 peer reviewed publications on on aging, dementia, Alzheimer’s and diabetes GNQA darnell2024biomedical
UMLS bodenreider2004umls Knowledge Base 2 million entities for 900K biomedical concepts BiomedRAG li2025biomedrag, CLEAR lopez2025clinical, Guo Guo2024, CliniqIR Abdullahi2024
ICD-10 cms2023icd10pcs Knowledge Base International statistical classification of diseases and related health problems Puts et al. puts2025icd10
HPO jsl2024hpo Knowledge Base Human phenotype ontology Albayrak et al. albayrak2024hpo, RAG-HPO garcia2024deepphenotyping, Kainer kainer2025effectiveness
MeSH Taxonomy Hierarchical organization of biomedical and health-related topics BMRETRIEVER xu2024bmretriever
AlzKB Matsumoto2025 Knowledge Graph Knowledge graph for Alzheimer’s disease research ESCARGOT Matsumoto2025, KRAGEN Matsumoto2024
DrugBank Wishart2006 Database Comprehensive database integrating detailed drug data with drug target information BMRETRIEVER xu2024bmretriever, Kim Kim2024
SPOKE morris2023spoke Knowledge Graph Integrating more than 40 publicly available biomedical knowledge sources of separate domains KG-RAG Soman2024
COD luo2024retrieval Knowledge Base Comprehensive ophthalmic dataset ChatZOC luo2024retrieval
HEALIE KGKakalou2024 Knowledge Graph Drawing from various medical ontologies, resources, and insights from domain experts KAKALOU et al. Kakalou2024
ADRD Caregiving parmanto2024calm Corpus Collection of resources supporting Alzheimer’s disease caregivers CaLM parmanto2024calm
StatPearls Bashir2024 Clinical Guide Collection of 9,330 clinical decision support articles available through NCBI Bookshelf MEDRAG xiong-etal-2024-benchmarking, i-MedRAG xiong2025iterative
Medical Textbooks jin2020disease Educational 18 core medical textbooks commonly used in USMLE preparation MEDRAG xiong-etal-2024-benchmarking, Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, i-MedRAG xiong2025iterative
Ophthalmology textbooks chen2024eyegpt Educational 14 specialized ophthalmology textbooks EyeGPT chen2024eyegpt
CheXpert irvin2019chexpert Image-Report 224,316 chest radiographs with associated reports CLEAR lopez2025clinical
Medical Task Datasets
Medical task datasets represent important benchmarking instruments for evaluating RAG systems within healthcare applications he2023survey, liu2024survey. These datasets simulate real clinical information challenges, providing structured frameworks to assess how systems process, interpret, and generate medical content.
These evaluation datasets can be systematically categorized according to distinct information-handling processes in the biomedical domain: biomedical information extraction, entity recognition, question answering (QA), biomedical multiple-choice examination, dialogue, and generation tasks. Table LABEL:tab:biomedical-comprehensive presents a comprehensive overview of the evaluation datasets employed across the surveyed biomedical RAG systems. Apart from works that make use of public biomedical datasets, several research groups Fukushima2025, steybe2025evaluation, wang2024diabetesrag in favor of human-curated questions, designing proprietary data for evaluation, and inviting healthcare professionals (e.g., physicians or nurses) to conduct assessments.
Table 9: Comprehensive Biomedical Tasks Datasets
Dataset Year Task Studies
ChemProt taboureau2010chemprot 2010 Information Extraction BiomedRAG li2025biomedrag, ChemProt
DDI segura2013semeval 2013 Information Extraction BiomedRAG li2025biomedrag, ChemProt
NCBI dogan2014ncbi 2014 Entity Recognition GeneGPT jin2024genegpt, NetMe DiMaria2024, RT li2024rt
BioASQ Tsatsaronis2015 2015 Information Retrieval BMRETRIEVER xu2024bmretriever, SeRTS Hu2024
MIMIC-III johnson2016mimic 2016 Text Summarization DR.KNOWS Gao2025, CliniqIR Abdullahi2024
BC5CDR li2016biocreative 2016 Entity Recognition RT li2024rt
DC3 eickhoff2019dc3 2019 Diagnostic classification CliniqIR Abdullahi2024
MIMIC-CXR johnson2019mimic 2019 Text Summarization Ascle Ascle
MedQuAD BenAbacha-BMC-2019 2019 QA Ascle Ascle
PubMedQA jin2019pubmedqa 2019 Multiple-choice BMRETRIEVER xu2024bmretriever, Liévin et al. Lievin2024, PodGPT jia2024podgpt
MMLU hendryckstest2021 2021 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, i-MedRAG xiong2025iterative, InfectA-Chat selcuk2025infecta, PodGPT jia2024podgpt
MedQA jin2020disease 2021 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, Liévin et al. Lievin2024, i-MedRAG xiong2025iterative, PodGPT jia2024podgpt
HealthSearchQA singhal2022large 2022 QA
MedMCQA pal2022medmcqa 2022 Multiple-choice Self-BioRAG jeong2024improving, JMLR Wang2024JMLR, Ascle Ascle, Liévin et al. Lievin2024, EyeGPT chen2024eyegpt, PodGPT jia2024podgpt
ChatDoctor li2023chatdoctor 2023 Dialogue EyeGPT chen2024eyegpt
MedAlpaca han2023medalpaca 2023 Dialogue EyeGPT chen2024eyegpt
GeneTuring hou2023geneturing 2023 Genomics QA GeneGPT jin2024genegpt
GIT li2023petailor 2023 Information Extraction BiomedRAG li2025biomedrag
MIRAGE xiong-etal-2024-benchmarking 2024 Multiple-choice & QA MEDRAG xiong-etal-2024-benchmarking
MedExpQA ALONSO2024102938 2024 Multilingual Medical QA MedExpQA ALONSO2024102938, PodGPT jia2024podgpt
PubMedRS-200 Lozano2024ClinfoAI 2024 Medical QA Clinifo.ai Lozano2024ClinfoAI
MedicineQA huang2024toolcalling 2024 Multi-round dialogue RagPULSE huang2024toolcalling
CELLS Guo2024 2024 Lay language generation Guo Guo2024
Evaluation Metrics
The biomedical RAG systems reviewed in this survey typically follow a three-stage pipeline, consisting of (1) retrieval, (2) reranking, and (3) generation. Consequently, the evaluation methodologies used across the surveyed studies can also be naturally organized according to these stages. Because each stage involves a distinct functional component of the RAG architecture, the choice of evaluation metrics naturally varies across stages. Moreover, as shown in Table LABEL:tab:biomedical-comprehensive, biomedical RAG systems are applied to a wide range of downstream tasks—such as NER, medical QA, dialogue, and clinical summarization—and the selection of evaluation metrics further depends on the specific characteristics of these tasks. In what follows, we review the evaluation metrics used in biomedical RAG research by following this pipeline-based structure.
Metrics for Retrieval
The retrieval stage evaluates how effectively a system surfaces and orders clinically relevant documents from large-scale biomedical corpora, with emphasis on ranking quality, relevance precision, and coverage of essential evidence. More broadly, evaluation in this stage follows an information-retrieval paradigm, where the goal is to quantify how well the retrieved set and its ordering align with the ground-truth relevance distribution. We review the core retrieval metrics commonly adopted in biomedical RAG systems—namely, Precision@k, Recall@k, Mean Average Precision (MAP@k), and normalized Discounted Cumulative Gain (nDCG@k)—and provide their formal definitions.
  1. (1)
    Precision@k
    $\text{Precision@k}=\frac{\sum_{i=1}^{k}\mathbf{1}[\text{doc}_{i}\text{ is relevant}]}{k}$ (7)
  2. (2)
    Recall@k
    $\text{Recall@k}=\frac{\sum_{i=1}^{k}\mathbf{1}[\text{doc}_{i}\text{ is relevant}]}{\lvert\text{Relevant Documents}\rvert}$ (8)
  3. (3)
    Mean Average Precision (MAP@k)
    $\text{AP@k}=\frac{1}{R}\sum_{i=1}^{k}\text{Precision@i}\cdot\mathbf{1}[\text{doc}_{i}\text{ is relevant}]$ (9)
    $\text{MAP@k}=\frac{1}{N}\sum_{j=1}^{N}\text{AP@k}(j)$ (10)
  4. (4)
    Normalized Discounted Cumulative Gain (nDCG@k)
    $\text{DCG@k}=\sum_{i=1}^{k}\frac{2^{rel_{i}}-1}{\log_{2}(i+1)}$ (11)
    $\text{nDCG@k}=\frac{\text{DCG@k}}{\text{IDCG@k}}$ (12)
Metrics for Reranking
The reranking stage refines the ordering of retrieved documents using more expressive scoring mechanisms (e.g., cross-encoder relevance models). Because its objective remains aligned with retrieval—prioritizing the most clinically relevant evidence—biomedical RAG systems typically evaluate reranking with the same IR metrics: Precision@k, Recall@k, MAP@k, and nDCG@k.Notably, in RAG applications, reranking quality is particularly significant because even small improvements in the top-ranked documents can substantially enhance the factual grounding and reliability of the final generated output.
Metrics for Generation
The generation stage is evaluated primarily through task-specific criteria, as the correctness and quality of the final output depend heavily on the nature of the downstream biomedical task. Because different applications—such as medical QA, clinical summarization, diagnostic classification, and dialogue—impose distinct requirements, the choice of evaluation metrics must be aligned with the characteristics of each task. In what follows, we group representative evaluation metrics into four categories: (1) generation metrics; (2) classification metrics; (3) named entity recognition metrics for biomedical IE tasks; and (4) dialogue metrics for assessing multi-turn clinical conversations.
Generation Metrics assess the quality of free-form textual outputs by measuring aspects such as content correctness, semantic alignment, and linguistic coherence. They are applicable to downstream biomedical generation tasks that require producing narrative responses, including medical QA, clinical report generation, and clinical summarization.
  1. (1)
    ROUGE-N
    $\text{ROUGE-N}=\frac{\text{Overlap of N-grams}}{\text{Total N-grams in reference}}$ (13)
    ROUGE-N measures the n-gram overlap between generated text and the reference text, commonly used for summarization and long-form generation.
  2. (2)
    ROUGE-L
    $\text{ROUGE-L}=\frac{\text{LCS}(X,Y)}{\text{Length of reference}}$ (14)
    ROUGE-L is based on the Longest Common Subsequence (LCS), capturing sentence-level structural similarity beyond exact n-gram matching.
  3. (3)
    BLEU
    $\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{4}w_{n}\log p_{n}\right)$ (15)
    BLEU evaluates the precision of n-grams and is widely adopted for machine translation and other structured text generation tasks.
  4. (4)
    BERTScore
    BERTScorebert-score evaluates semantic similarity between a generated text $X$ and a reference text $Y$ by aligning their token embeddings from a pretrained language model. Let $e_{x}$ and $e_{y}$ denote contextual embeddings of tokens $x\in X$ and $y\in Y$, respectively. The precision $P$ and recall $R$ are defined as
    $P=\frac{1}{|X|}\sum_{x\in X}\max_{y\in Y}\cos\bigl(e_{x},e_{y}\bigr),\quad R=\frac{1}{|Y|}\sum_{y\in Y}\max_{x\in X}\cos\bigl(e_{y},e_{x}\bigr),$ (16)
    and the final BERTScore is computed as the $F_{1}$ score:
    $\text{BERTScore}_{F_{1}}=\frac{2\cdot P\cdot R}{P+R}.$ (17)
    Compared to lexical overlap metrics, BERTScore offers a more contextualized evaluation by operating in the embedding space.
  5. (5)
    BLEURT
    BLEURTsellam-etal-2020-bleurt is a learned evaluation metric based on pretrained language models fine-tuned on human-rated text pairs. Given a generated text $X$ and reference text $Y$, BLEURT encodes both sequences with a transformer encoder and predicts a scalar quality score:
    $\text{BLEURT}(X,Y)=f_{\theta}\bigl(\text{Enc}(X),\,\text{Enc}(Y)\bigr),$ (18)
    where $\text{Enc}(\cdot)$ denotes contextual embeddings and $f_{\theta}$ is a regression head trained to approximate human judgments. BLEURT captures semantic adequacy and fluency beyond surface-level overlap.
  6. (6)
    RaTEScore
    RaTEScorezhao2024ratescore is an entity-based metric for radiology report generation. It extracts medical entities and their relations from both generated and reference reports, matches them with synonym awareness, weights them by clinical importance, and computes an F1-style score. This makes the metric focus on factual, clinically meaningful consistency rather than surface wording, achieving higher alignment with expert judgment.
Classification metrics evaluate the correctness of discrete predictions made by a model, such as assigning labels or choosing among predefined options. These metrics are well-suited for tasks like diagnosis classification and multiple-choice question answering.
  1. (1)
    Accuracy (ACC) measures the proportion of correctly predicted instances:
    $\text{ACC}=\frac{\text{\# correct predictions}}{\text{\# total instances}}$ (19)
  2. (2)
    Precision, Recall, and F1 for a single class are defined as:
    $\text{Precision}=\frac{TP}{TP+FP},\quad\text{Recall}=\frac{TP}{TP+FN},\quad\text{F1}=\frac{2\cdot\text{Precision}\cdot\text{Recall}}{\text{Precision}+\text{Recall}}$ (20)
  3. (3)
    Macro-F1 is commonly used in multi-class and multi-label classification settings. It computes the F1 score independently for each class and then averages them, giving equal weight to both frequent and rare classes:
    $\text{Macro-F1}=\frac{1}{C}\sum_{c=1}^{C}\frac{2\cdot\text{Precision}_{c}\cdot\text{Recall}_{c}}{\text{Precision}_{c}+\text{Recall}_{c}}$ (21)
  4. (4)
    Micro-F1 is another metric for multi-class and multi-label classification tasks. It aggregates true positives, false positives, and false negatives across all classes before computing a global F1 score, thus favoring frequent classes:
    $\text{Micro-F1}=\frac{2\cdot\sum_{c=1}^{C}TP_{c}}{2\cdot\sum_{c=1}^{C}TP_{c}+\sum_{c=1}^{C}FP_{c}+\sum_{c=1}^{C}FN_{c}}$ (22)
AUROC and AUPRC are probability-based evaluation metrics commonly used to assess the performance of binary classifiers across different decision thresholds. We notice that they are rarely reported in the surveyed works, as LLMs typically generate free-text outputs or are prompted to select a single label from a predefined set rather than produce explicit class probabilities.
Named Entity Recognition (NER) metrics measures how accurately a system identifies and labels biomedical entities in text, and are widely used for assessing biomedical information extraction tasks.
  1. (1)
    Token-level F1 evaluates NER performance at the token granularity by computing the F1 score over token-wise precision and recall:
    $\text{Token-F1}=\frac{2\cdot P\cdot R}{P+R}$ (23)
  2. (2)
    Entity-level F1 assesses how accurately complete entity spans are identified, considering correct entities, boundary errors, and missed entities:
    $\text{Entity-F1}=\frac{2\cdot\lvert\text{Correct Entities}\rvert}{2\cdot\lvert\text{Correct Entities}\rvert+\lvert\text{Wrong Boundaries}\rvert+\lvert\text{Missed Entities}\rvert}$ (24)
Medical Dialogue evaluation is inherently complex, as it must account for both the linguistic quality of generated responses and the effectiveness of completing dialogue-level clinical tasks. In practice, evaluation combines automatic metrics with human or LLM-based judgments, since purely automated measures are often insufficient for safety-sensitive biomedical scenarios.
  1. (1)
    Generation quality (automatic): Metrics such as ROUGE, BLEU, BERTScore, and DialogRPTgao-etal-2020-dialogue are used to automatically assess lexical overlap, semantic similarity, and response coherence. These metrics provide scalable proxies for linguistic quality but do not capture clinical correctness.
  2. (2)
    Task success (primarily automatic): Indicators such as task completion rateli2021how and slot accuracyaksu2024granular measure whether the system achieves predefined clinical or informational goals. These metrics can be computed automatically when gold annotations are available, although human assessment is sometimes required for semantically defined success criteria.
  3. (3)
    Safety and factuality (human/LLM-based): Evaluation of hallucination, factual consistency, and clinical appropriateness typically relies on expert reviewers or medically aligned LLM judges Croxford2025, metric-survey, as automated approaches have limited ability to detect subtle or domain-specific medical errors. Automatic hallucination or consistency scores may be used as auxiliary signals, but do not replace human or expert-guided evaluation.
Together, these complementary evaluation strategies constitute a diverse assessment framework, and surveyed studies typically adopt one or more of these metrics to evaluate the quality of RAG-based medical dialogue systems.
Applications in biomedicine
The application of RAG in the biomedical domain represents a significant advancement in healthcare informatics, combining the strengths of LLMs with domain-specific knowledge retrieval. This integration helps mitigate challenges inherent to healthcare applications—factual accuracy, domain specificity, knowledge recency, and explainability—that conventional generative AI approaches struggle to overcome liu2025improving. In this section, we analyze 5 key domains in which we identified that biomedical RAG systems have demonstrated substantial utility: clinical decision support, medical documentation, precision medicine, education, and clinical research.
Clinical Decision Support Systems
Medical Question Answering
Medical question-answering systems provide clinicians and patients with access to accurate healthcare information at the point of need Soman2024. This application is particularly valuable because it enables rapid retrieval of evidence-based information, supporting clinical decision-making while reducing the cognitive load on healthcare providers and improving information access for patients with varying levels of health literacy.
Recent advancements have produced several novel medical QA systems with distinctive approaches to knowledge retrieval. Clinfo.ai Lozano2024ClinfoAI provides evidence-based healthcare QA by retrieving information from trusted sources like PubMed using the Entrez API, ensuring responses reflect current medical evidence. RALL Guo2024 employs Dense Passage Retriever to generate lay language explanations of medical concepts, addressing the growing need for accessible medical information. Self-BioRAG jeong2024improving implements the MedCPT retriever with self-reflection techniques to improve accuracy for complex clinical questions with multiple components, while WeiseEule Aftab2024 evaluates multiple retrievers (MedCPT, BioBERT, BioGPT) with Pinecone DB to handle diverse medical information needs across specialties. These medical QA systems collectively demonstrate RAG’s potential for healthcare information access, providing contextually relevant, evidence-based responses to complex medical queries for both clinical decision support and patient education.
Diagnostic and Treatment Decision Support
Diagnostic and treatment decision support systems assist clinicians in navigating complex clinical pathways through evidence retrieval and synthesis Sutton2020CDSS. They can help healthcare providers identify relevant evidence, recognize patterns, and identify appropriate interventions, ultimately improving diagnostic accuracy and treatment selection while reducing cognitive biases and diagnostic errors.
Several innovative implementations showcase RAG’s potential in clinical decision support across medical specialties. Expert-Guided LLMs Lammert2024ExpertLLM support specialized decision-making in precision oncology by retrieving relevant genomic information and matching it with treatment guidelines and clinical trial data. DR.KNOWS Gao2025 facilitates diagnostic prediction by connecting EHRs with medical knowledge repositories through knowledge graph-enhanced retrieval that models complex relationships between symptoms, diseases, and treatments. CliniqIR Abdullahi2024 combines MedCPT and BM25 retrieval to match patient symptoms with potential diagnoses, while SurgeryLLM ong2024surgeryllm delivers surgical decision support by retrieving relevant procedures, techniques, and complication data. RECTIFIER unlu2024retrieval employs OpenAI Text-Embedding with Faiss for clinical trial screening, helping match heart failure patients with appropriate clinical trials by analyzing eligibility criteria against patient data. By connecting patient-specific data with evidence-based medical knowledge, these RAG systems help clinical decision-making while reducing the cognitive burden on healthcare providers, potentially improving both efficiency and quality of care.
Rare Disease Identification and Management
Rare disease identification and management represents a particularly challenging healthcare domain due to limited literature, dispersed expertise, and complex symptom patterns MartinezDeMiguel2022RareDis. RAG systems are especially valuable in this context because they can aggregate and synthesize knowledge from diverse sources, helping clinicians recognize unusual patterns and connect them with appropriate diagnostic and management strategies for conditions they may encounter only rarely in clinical practice.
Recent research has produced several specialized rare disease applications leveraging different knowledge retrieval approaches. Zelin et al. Zelin2024RareDisease developed a RAG system specifically focused on rare disease diagnosis, connecting uncommon symptom patterns with rare disease knowledge drawn from specialized databases. RAG-HPO garcia2024deepphenotyping utilizes FastEmbed for automated deep phenotyping of rare genetic disorders, linking phenotypic observations with potential genetic causes through human phenotype ontology mapping. Yang et al. Yang2024RDguru created RDguru, an intelligent conversational agent that employs a multi-source fusion diagnostic model combining GPT-4, PheLR, and phenotype matching strategies to significantly improve diagnostic recall for rare diseases. A recent study kim2024phenotype has also evaluated the capabilities of LLMs and RAG for phenotype-driven gene prioritization in rare genetic disease diagnosis. These applications demonstrate how RAG systems can help bridge knowledge gaps in rare disease management by connecting dispersed information to support timely diagnosis and appropriate treatment strategies for conditions where clinical expertise is often limited.
Clinical Report Generation
Automatic and semi-automatic report generation can help healthcare providers by significantly reducing the time they spend on documentation, allowing them to dedicate more time to patient care sloan2024automated. A critical area of report generation is radiology, which involves image analysis, document consultation, and data evaluation hamamci2024ct2rep. This process faces multiple challenges including ensuring factual accuracy, maintaining clinical relevance, minimizing hallucinations, and providing sufficient interpretability for clinical adoption alfarghaly2021automated.
RAG frameworks have been proposed to address these challenges by grounding language generation in relevant retrieved content, resulting in more accurate and trustworthy outputs xiong2024benchmarking. Recent works have demonstrated the effectiveness of RAG in this domain. For example, multimodally-aligned embeddings with general-domain LLMs such as GPT-4 ranjit2023retrieval, ensuring relevant content retrieval and mitigating hallucinations while allowing customization based on clinical intent. To enhance interpretability, some authors alam2024towards proposed a multi-agent RAG framework utilizing concept bottleneck models that guide report generation with clinically meaningful concept vectors, bridging the gap between performance and explainability. RAG pipelines bernardi2024report have also been used to improve image-text matching techniques, incorporating semantic segmentation and triple contrastive loss, leading to more accurate and fluent reports with reduced errors. Differently, LaB-RAG song2024lab employs label-boosted retrieval by converting medical images into descriptive text labels, which are then used for text-based retrieval and report generation, bypassing the need for direct image processing by the LLM. Lastly, some authors sun2024fact have proposed fact-aware multimodal retrieval strategies to address factual inconsistencies in generated reports by leveraging RadGraph-based factual mining and multimodal retrievers, ensuring high-quality reference retrieval and improved factual completeness.
Precision Medicine Applications
Genomic Medicine
Genomic medicine involves analyzing genetic data to inform clinical decision-making. The rapid knowledge evolution and inherent complexity in variant interpretation within genomic medicine impose significant cognitive burdens on clinicians, particularly given the exponential expansion of genomic information and the challenges in ascertaining variant pathogenicity burton2012genomic.
Several innovative RAG applications address these challenges in genomic medicine through specialized knowledge retrieval. FAVOR-GPT li2024favorgpt utilizes OpenAI Text-Embedding with Weaviate DB to facilitate genome variant annotations, connecting genetic variants with clinical significance and supporting evidence from multiple databases. Lu et al. Lu2025BoostingGPT developed a genomics analysis system using Azure AI Search with keyword matching for more accurate genetic variant interpretation within clinical contexts. RAG-HPO garcia2024deepphenotyping connects clinical observations with genetic variants in rare disorders through sophisticated knowledge mapping, while JGCLLM Fukushima2025 implements GLuCoSE-base-ja vectors to provide genetic counseling support that translates complex genomic findings into clinically meaningful information. By connecting genomic findings with clinical knowledge through appropriate retrieval mechanisms, these RAG systems support the clinical use of genetic information for tasks such as disease inference Lu2025BoostingGPT as well as phenotypic and genotypic analysis garcia2024deepphenotyping, thereby enabling more personalized treatment decisions in the emerging era of genomic medicine.
Personalized Health Management
Personalized health management, a tailored healthcare approach, involves generating individualized health information and recommendations based on patient-specific characteristics, preferences, and health status Quazi2022HCI. This application mitigates a important healthcare challenge: the need for more effective and patient-centered care that optimizes treatment adherence, self-management, and health outcomes, in contrast to traditional standardized approaches that often fail to accommodate specific patient needs, literacy levels, and medical conditions.
Recent research has produced several personalized health management systems employing different knowledge retrieval strategies. HEALIE Kakalou2024 leverages knowledge graphs to generate personalized medical content incorporating both clinical knowledge and socio-economic factors relevant to individual patients. RISE wang2024diabetesrag implements OpenAI Text-Embedding with FAISS to provide personalized responses to diabetes-related inquiries, helping patients navigate complex self-management requirements. A Dual RAG System Lee2024 combines dense retrieval and BM25 with language-specific tokenizers to support diabetes management across multiple languages, expanding access to personalized health information for diverse populations. RAP azimi2025evaluation utilizes Amazon Titan Text Embeddings to provide personalized nutrition-related question answering, supporting dietary management for various health conditions. These personalized health management systems demonstrate RAG’s ability to bridge the gap between complex medical information and individual patient needs through contextually relevant knowledge retrieval, ultimately supporting better self-management and improved health outcomes through personalized guidance.
Healthcare Education
Medical education and training systems support the development of healthcare professionals through evidence-based learning resources and simulated clinical experiences Lemmetty2006HIS. These systems help address workforce needs by enabling more efficient knowledge acquisition, supporting the development of clinical reasoning skills, and allowing for personalized learning experiences that adapt to individual learner needs and knowledge gaps.
ChatENT long2024chatent leverages Retrieval-Augmented Language Modeling (RALM) to support medical education in Otolaryngology-Head and Neck Surgery (OHNS). EyeGPT chen2024eyegpt transforms a general LLM into an ophthalmology-specialized assistant through prompt engineering, fine-tuning with eye-specific terminology, and RAG from ophthalmology textbooks, supporting both medical education and patient inquiries. Sevgi et al. sevgi2024medical developed targeted ophthalmology applications, including EyeTeacher for generating educational questions and EyeAssistant for clinical support. Both systems are enhanced by integrating clinical guidelines and peer-reviewed documents as retrieval sources. In nursing education, Zhao et al ZHAO2025104284 leveraged RAG and semantic search technology to answer questions from the Chinese Nursing Licensing Exam, showing the potential for specialized knowledge retrieval in professional certification. By connecting learning objectives with relevant medical knowledge through appropriate retrieval mechanisms, these RAG-based education systems help support healthcare practice and skill training.
Clinical Research
Clinical research application tools are designed to support clinical studies by enhancing efficiency, leveraging knowledge retrieval and synthesis to accelerate the development of medical knowledge and therapeutic innovations. ge2024leveraging.
Recent literature highlights several innovative RAG applications in clinical research leveraging different retrieval approaches. PodGPT jia2024podgpt implements the BAAI embedding model for scientific literature retrieval informing research design and methodology development. LITURAt Peasley2025 utilizes the Entrez API for PubMed retrieval, supporting scientific data analysis in clinical research through relevant literature identification. These applications showcase RAG’s ability to optimize and improve the clinical research process through intelligent knowledge retrieval, highlighting its potential as a valuable tool for both research and education, and supporting the translation of research findings into clinical practice.
Discussion
In this section, we summarize the key observations drawn from technological approaches, datasets, evaluation resources, and biomedical applications, examines the limitations of existing work, and further discusses the major challenges faced by RAG systems in the biomedicine domain while outlining promising directions for future research and development.
Summary of technological approaches
In Section.III, we reviewed three key technological foundations of biomedical RAG systems identified in our survey: retrieval, reranking, and generation. Retrieval methods span traditional sparse models such as BM25bm25, multiple classes of dense retrievers—including general PLMs, commercial embedding APIs, domain-adapted biomedical models, LLM-based embeddings, and specialized architectures such as MedCPTjin2023medcpt and BMRetrieverxu2024bmretriever, and multimodal frameworks. Knowledge-graph–augmented retrieval further enhances precision by leveraging structured biomedical ontologies and explicit entity–relation pathways. Benefiting from its efficiency and proven generalizability, BM25bm25 remains the leading approach for retrieval; however, it is limited to textual resources. Emerging image-based and multimodal systems remain limited, partly because single-vector indexingXia2024RULE, pmlr-v219-ranjit23a, Sun2024FactAware—widely used in text retrieval—may underrepresent the heterogeneous characteristics of multimodal data, making their indexing and retrieval methodologies for multimodal RAG still worthy of further exploration.
Reranking techniques provide a second-stage selection mechanism, aiming to enhance the quality of the initial rank with more advanced and algorithmically complex strategies. Existing systems employ statistical approaches (e.g., RRF, MMR), transformer-basedjia2024podgpt, adejumo2024retrieval and cross-encoder modelsjin2023medcpt, LLM-based relevance classifiersdarnell2024biomedical, craig2024lmrac, as well as domain-specific rerankers designed to better capture biomedical semantics. The main challenge for reranking models is generalization at inference time, as they often rely on annotated samples. Creating application-specific rank–gold labels for existing datasets is expensive, while automatic annotation carries risks of reduced accuracy, forcing developers to carefully balance this trade-off in practice.
Lastly, on the generation side, both open-source and commercial LLMs are used extensively, alongside an increasing number of biomedical-specialized models (see Table LABEL:tab:llm_taxonomy), to generate content benefiting from the retrieved biomedical resources. We also find that domain adaptation—whether applied to retrievers, rerankers, or generators—enhances terminology handling, conceptual grounding, and factual reliability. Building on these observations, commercial LLMs are easy to use and can deliver competitive performance, whereas open-source models—especially biomedical-specialized ones—allow continued training or fine-tuning for specific tasks and even modifications to the model architecture, enabling more flexible application across diverse tasks. However, this flexibility also imposes additional computational resource demands on developers.
For the future development of RAG, we believe that continued progress in multimodal retrieval and knowledge graph integration will expand the scope of information accessible to RAG systems and improve their ability to generate contextually informed, clinically aligned outputs. This underscores the growing importance of cross-modal fusion strategies and targeted domain adaptation in advancing biomedical RAG capabilities.
Summary of dataset and evaluation approaches
In Sec.IV, we reviewed datasets relevant to the development and evaluation of biomedical RAG systems, categorizing them into Knowledge Source Datasets (Table LABEL:tab:med_knowledge) and Medical Task Datasets (Table LABEL:tab:biomedical-comprehensive). Biomedical knowledge sources are highly diverse, and this heterogeneity poses significant challenges for retrieval, as different sources often exhibit distinct linguistic styles and vocabulary patterns.
On the evaluation side, current benchmarks mainly rely on various Medical Task Datasets, with a particular focus on knowledge-intensive tasks such as Multiple-Choice Examinations (e.g., MedQA jin2020disease, MedMCQA pal2022medmcqa), which simulate medical licensing assessments, and a range of Question Answering (QA) tasksBenAbacha-BMC-2019, singhal2022large, hou2023geneturing, Lozano2024ClinfoAI, ALONSO2024102938. However, existing evaluations still lack tasks that are more concrete and aligned with real-world clinical scenarios. We also observe a strong dominance of English-language datasets and a heavy reliance on textual benchmarks. Substantial future efforts are needed to build richer cross-lingual and multi-modal evaluation resources that better reflect the diversity and complexity of biomedical practice.
For evaluation metrics, we follow the retrieval–reranking–generation pipeline and review the metrics used at each stage accordingly. Retrieval and reranking are typically assessed using standard information-retrieval measures such as recall, precision, nDCG, and MRR, while generation metrics vary widely depending on the specific downstream task. To provide a clearer overview, we group generation-level evaluations into four major task types, each associated with its own metric family. Although many automatic metrics exist, they often struggle to capture the nuanced reasoning quality required in complex biomedical dialogue and decision-support scenarios metric-survey. As a result, human evaluation remains essential. However, due to the substantial cost of expert annotation, LLM-based evaluators are increasingly adopted Croxford2025, metric-survey. These models can follow instructions and emulate human evaluation criteria, thereby offering a more scalable alternative to manual assessment, although they may also introduce risks such as hallucination and evaluation bias. Therefore, developing evaluation frameworks that are both low-cost and highly accurate remains an important direction for future research.
Summary of RAG applications in biomedicine
Biomedical RAG applications (Sec. IV) span a wide range of healthcare domains—from clinical decision support and precision medicine to medical education and clinical research—highlighting RAG’s capacity to enhance factual grounding, interpretability, and domain specificity for AI-driven healthcare solutions. Across the surveyed literature, RAG-based systems have demonstrated utility in various medical specialties, including oncologyLammert2024ExpertLLM, radiologysloan2024automated, ophthalmologychen2024eyegpt, geneticsli2024favorgpt, Lu2025BoostingGPT, Fukushima2025, nursingZHAO2025104284, and rare disease diagnosisZHAO2025104284, Zelin2024RareDisease, garcia2024deepphenotyping. These domain-specific implementations suggest that retrieval-augmented approaches can help connect LLMs with biomedical evidence, thereby supporting grounded question answering, enhancing the consistency of clinical reasoning processes, and contributing to clearer and more interpretable outputs for both patient-facing and clinician-facing use cases. However, the surveyed applications also reveal several practical constraints. Retrieval steps often introduce latency and increase token length, which can reduce system efficiency and limit scalability in real clinical workflowsAmmar2025. In addition, current retrievers still struggle with fine-grained biomedical semantics—such as distinguishing closely related clinical entities or integrating cross-modal evidence—which can lead to suboptimal grounding even when high-quality sources existTan2024. Addressing these issues will be important for deploying RAG systems in settings that demand both timely responses and high diagnostic precision.
Challenges and Future Direction
Trustworthiness
Recent research indicates that RAG systems can be both maliciously and unintentionally manipulated to produce unreliable decisions and potentially harm humans Fan2024RAG, deng2024pandora. Such vulnerabilities are particularly unacceptable in the biomedical domain due to the potential for severe consequences.
Several dimensions of trustworthiness in biomedical RAG systems raise critical concerns. Robustness refers to the system’s ability to exclude malicious content introduced by attackers or filter inadvertent misinformation contained in web corpora that might contaminate the knowledge source gao2023retrieval. Although certain re-ranking methodologies described in Section 3.2 can partially mitigate these issues, significant challenges remain in biomedical RAG applications. Explainability requires a comprehensive understanding of user intent to generate reliable responses aligned with ground-truth documents. A notable challenge emerges from diverse linguistic styles in input information, creating misalignment between user queries and knowledge sources. For instance, in medical QA systems, users may formulate queries in conversational language while knowledge sources are typically based on academic publications lee2024enhancing. This linguistic discrepancy poses substantial challenges for biomedical information retrieval, consequently diminishing the performance of biomedical RAG systems. Fairness dictates that trustworthy RAG systems must avoid discriminatory outputs during the generation process. Due to the uneven distribution of demographic attributes in training corpora he2023survey, which may be present in both the pre-training data of LLMs and the knowledge sources utilized in RAG systems, biomedical RAG applications are susceptible to intrinsic biases. Addressing the fairness concerns also represents an important direction for future research.
Multilinguality
Among our surveyed publications, English remains the dominant language in all Biomedical RAG systems. This predominance is primarily attributed to the dominance of English in public biomedical knowledge repositories and medical task datasets. The performance of biomedical RAG systems in other languages, particularly those with limited resources, remains inadequately explored neveol2018clinical, Englishllm. This fact presents two significant challenges in the development of multi-lingual biomedical RAG systems.
First, training LLMs specialized in low-resource languages is inherently difficult, which diminishes the generative capabilities of Biomedical RAG in other linguistic contexts. Second, the scarcity of available knowledge sources in non-dominant languages significantly constrains the augmentation potential afforded by knowledge retrieval mechanisms. Although some researchers have attempted to address this limitation through synthetic data generated by LLMs or translated from resource-rich languages mondshine2025beyond, consistent and sustained efforts are still required to overcome this challenge in multilingual biomedical information retrieval and generation.
Multimodality
Another challenge lies in extending retrieval capabilities beyond textual data to encompass multi-modalities highly prevalent in the biomedical domain, including images (e.g. X-ray johnson2019mimic, pelka2018roco, lin2023pmc), time series (e.g. ECG pmlr-v225-yu23b), sound recordings (e.g. lung sounds recorded fraiwan2021dataset), and video gupta2023dataset. Among the studies we identified, the overwhelming majority focus exclusively on text-based retrieval, with only a limited subset addressing image-based retrieval, while other modalities remain largely unexplored. Multi-modal data presents several intractable challenges for retrieval systems. For instance, in medical imaging, although we can utilize a visual encoder model as a retriever, high intra-class variance significantly complicates image retrieval processes zhong2024class. Specifically, identical medical conditions can manifest with remarkable heterogeneity across visual representations; for instance, images of histologically identical tumors may exhibit substantially different morphological characteristics when visualized across different patients or at varying stages of disease progression. Attention also should be paid to developing cross-modal RAG in the biomedical domain and improving its accuracy and trustworthiness in the future.
Restricted Computing Resources
The application of RAG in the biomedical domain faces significant limitations due to computing resource constraints, which restrict its utility in practical scenarios. Within the RAG framework, LLMs play a crucial role in contextual understanding and human-like response generation based on retrieved knowledge Fan2024RAG. However, these models typically require high-performance computing clusters to operate effectively. For instance, the minimum deployment unit of DeepSeek V3’s decoding stage consists of 40 nodes with 320 GPUs deepseekv3technicalreport.
Empirical research consistently demonstrates that LLMs with larger parameter counts generally exhibit superior capacity for understanding and generation tasks brown2020language, hoffmann2022training. Nevertheless, in real-world healthcare environments such as hospitals, maintaining high-performance computing infrastructure is often impractical and cost-prohibitive. Several promising approaches have emerged to address these constraints. One solution involves developing smaller, distilled LLMs derived from their larger counterparts xu2024survey. Models such as QWQ qwen32b_2024 and Mistral Small 3.1 mistral_small_24b_2024 demonstrate competitive performance despite having significantly reduced parameter counts. Applying portable LLMs khan2024portllm as generation models is also a promising method. Another viable approach incorporates RAG systems that combine local knowledge bases with external LLM API services. This hybrid configuration presents an acceptable compromise when computing resources are limited, although it introduces potential privacy concerns, as local data must be transmitted to commercial providers for response generation.
Privacy
Privacy concerns in biomedical RAG systems exist throughout both retrieval and generation processes. During knowledge retrieval, reference documents frequently contain patients’ clinical histories, which may include personally identifiable information that individuals would prefer to keep confidential meystre2010automatic. Such information handling might not comply with stringent regulatory frameworks such as the General Data Protection Regulation GDPR2016 and the Health Insurance Portability and Accountability Act HIPAA1996. Furthermore, the generation phase presents additional challenges, particularly when utilizing commercial LLMs for biomedical RAG development. This approach necessitates the transmission of sensitive data across the internet, significantly increasing the risk of data breaches and unauthorized access. Consequently, developing biomedical RAG systems with robust privacy-preserving mechanisms remains a critical direction for future research endeavors.
Conclusion
In conclusion, this comprehensive survey has provided a detailed overview of the application of RAG in the biomedical domain. It has examined the basic framework of biomedical RAG and discussed its retriever, reranker, and generation modules, as well as explored KG-enhanced biomedical RAG and multi-modal biomedical RAG. The survey has also explored clinical applications where biomedical RAG can be applied and reviewed recent studies in this field. Additionally, the paper has identified key challenges, including trustworthiness, multilingual capabilities, multimodal integration, limited computational resources, and privacy concerns. Overall, this work provides researchers and practitioners with a thorough understanding of the current state of biomedical RAG systems, ultimately contributing to the development of more accurate, accessible, and impactful healthcare AI systems that can meaningfully enhance patient care through evidence-based information access and support.